### v1

# Huppert’s conjecture for finite simple exceptional groups of Lie type

Abstract.
Let $G$ be a finite group and let ${\mathrm{cd}}(G)$ be the set of all complex irreducible character degrees of $G.$ In this paper, we show that if ${\mathrm{cd}}(G)={\mathrm{cd}}(H),$ where $H$ is a finite simple exceptional group of Lie type, then $G\cong H\times A,$ where $A$ is an abelian group. This completes the verification of Huppert’s Conjecture for all finite simple exceptional groups of Lie type.
Key words and phrases:
character degrees; exceptional groups of Lie type; Huppert’s Conjecture
2000 Mathematics Subject Classification:
Primary 20C15, secondary 20D05, 20G40
1. Introduction
Let $G$ be a finite group. Denote by ${\mathrm{Irr}}(G)$ the set of all complex irreducible characters of $G$. Let ${\mathrm{cd}}(G)$ be the set of character degrees of $G$, that is, ${\mathrm{cd}}(G)=\{\chi(1):\chi\in{\mathrm{Irr}}(G)\}.$ One of the main questions in the character theory of finite groups is to determine which properties of a finite group can be determined by its character table. Note that the character degree set of a group can be read off from the character table. Thus it is interesting to ask how much the character degree set of a group knows about the group itself.
Apparently, we do not expect much since the character degree set of a finite group $G$ and the character degree set of a direct product $G\times A$, where $A$ is an abelian group, coincide. In fact, G. Navarro [21] showed that the character degree set alone cannot determine the solvability of the group. He constructed a finite perfect group $H$ and a finite solvable group $G$ such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. Furthermore, Navarro and Rizo [23] found a finite perfect group and a finite nilpotent group with the same character degree set.
On the other hand, B. Huppert proposed in [9] that all finite nonabelian simple groups are essentially determined by the set of their character degrees, that is, if $H$ is a finite nonabelian simple group and $G$ is a finite group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$, then $G\cong H\times A$ for some finite abelian group $A$. The main purpose of this paper is to prove the following theorem.
Theorem.
Let $G$ be a finite group and let ${H}$ be a finite simple exceptional group of Lie type. If ${\mathrm{cd}}(G)={\mathrm{cd}}({H}),$ then $G\cong H\times A,$ where $A$ is abelian.
Since the conjecture has been verified for all finite simple exceptional groups of Lie type of Lie rank at most 2, (see [9, 25, 26, 28, 29, 30]), we only need to verify the conjecture for the following finite simple groups: $\textrm{F}_{4}(q),~{}{}^{2}\textrm{E}_{6}(q),~{}\textrm{E}_{6}(q),~{}\textrm{E}_{7}(q),~{}\textrm{E}_{8}(q)$, where $q$ is a prime power.
We should mention that Huppert’s Conjecture has been verified for all sporadic simple groups (see [27] and the references therein), the alternating groups [1] and many finite simple classical groups of small Lie rank (see [11]). The conjecture remains widely open for finite simple classical groups of large rank.
In [12], we extend the conjecture to include certain quasi-simple groups. Recently, an analog of Hupper’s Conjecture was proposed by G. Qian in [14, Problem 20.79] using codegrees. Recall that the codegree of a character $\chi\in{\mathrm{Irr}}(G)$ is defined to be the ratio $|G:{\mathrm{ker}}(\chi)|/\chi(1)$, where ${\mathrm{ker}}(\chi)$ denotes the kernel of $\chi$. Let $\textrm{cod}(G)$ be the set of all codegrees of $G$. It is conjectured by G. Qian that if $G$ is a finite group and $H$ is a nonabelian simple group such that $\textrm{cod}(G)=\textrm{cod}(H)$, then $G\cong H.$ This conjecture has been verified for alternating groups, all sporadic simple groups, all finite simple exceptional groups of Lie type and finite simple projective special linear groups (see [24] and the references therein).
We now describe our approach to the proof of Huppert’s Conjecture. In [9], Huppert proposed a $5$-step strategy to verify his conjecture and this has been used successfully to verify the conjecture for many finite simple groups. In [1], we modified his method to verify the conjecture for alternating groups. We will make use of this method again together with the data from [16] to prove the main theorem.
Suppose that $G$ is a finite group and $H$ is a finite nonabelian simple group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. To verify Huppert’s Conjecture for the nonabelian simple group $H$, we need to prove the following.
Step 1: Show that $G^{\prime}=G^{\prime\prime}$;
Step 2: If $L/M$ is any nonabelian chief factor of $G$, then $L/M\cong H;$
Step 3: If $L$ is a finite perfect group and $M$ is a minimal normal elementary abelian subgroup of $L$ such that $L/M\cong H$, then some degree of $L$ divides no degree of $H$.
Step 4: If $T$ is any finite group with $H\unlhd T\leq\mathop{\rm Aut}\nolimits(H)$ and $T\neq H$, then ${\mathrm{cd}}(T)\nsubseteq{\mathrm{cd}}(H)$.
To verify Steps $1$ and $2$, we will follow the argument developed by Huppert. For Step 3, we first see that either ${\mathbf{C}}_{L}(M)=M$ or ${\mathbf{C}}_{L}(M)=L$. For the first possibility, we use a result due to Guralnick and Tiep [8] on the non-coprime $k(GV)$ problem. If the latter holds, then $L$ is a quasi-simple group. This and Step 4 can now be dealt with using the data from [16].
The paper is organized as follows. We collect some results needed to verify Steps 1 and 3 in Section 2. In Section 3, we present all the information about the character degrees of the finite simple exceptional groups of Lie type that we will need for the proof of the main theorem. We will verify Steps 1–4 in Sections 4–7, respectively.
Notation.
If $n$ is an integer then we denote by $\pi(n)$ the set of all prime divisors of $n.$ If $G$ is a group, we write $\pi(G)$ to denote the set of all prime divisors of the order of $G.$ Let $\rho(G)=\cup_{\chi\in{\mathrm{Irr}}(G)}\pi(\chi(1))$ be the set of all primes which divide some irreducible character degrees of $G.$ The largest character degree of $G$ will be denoted by $b(G).$ For a group $G,$ let $k(G)$ be the number of conjugacy classes of $G.$ If $N\unlhd G$ and $\theta\in{\mathrm{Irr}}(N),$ then the inertia group of $\theta$ in $G$ is denoted by $I_{G}(\theta).$ The set of all irreducible constituents of $\theta^{G}$ is denoted by ${\mathrm{Irr}}(G|\theta).$ Finally, we denote by $\Phi_{n}=\Phi_{n}(q),$ the cyclotomic polynomial in variable $q.$ Other notation is standard.
2. Preliminaries
We collect in this section various results that are needed for the proofs of the main theorem. We start with the following lemma which will be used to verify Steps $2$ and $3$
Lemma 2.1.
If $S$ is a nonabelian simple group, then there exists a nontrivial irreducible character $\theta$ of $S$ that extends to ${\mathop{\rm Aut}\nolimits}(S).$ Moreover the following holds:
$(i)$ If $S$ is an alternating group of degree at least $7,$ then $S$ has two characters of consecutive degrees $n(n-3)/2$ and $(n-1)(n-2)/2$ that both extend to ${\mathop{\rm Aut}\nolimits}(S).$
$(ii)$ If $S$ is a sporadic simple group or the Tits group, then $S$ has two nontrivial irreducible characters of coprime degrees which both extend to ${\mathop{\rm Aut}\nolimits}(S).$
$(iii)$ If $S$ is a simple group of Lie type then the Steinberg character ${\mathrm{St}}_{S}$ of $S$ of degree $|S|_{p}$ extends to ${\mathop{\rm Aut}\nolimits}(S).$
Proof.
This is essentially [2, Theorems 2,3,4]. However for sporadic simple groups and the Tits group, we will choose different pairs of coprime degrees. These two degrees are listed in Table 1. ∎
We also need the following number theoretic result. It will be used to show that certain character degree of the simple exceptional groups of Lie type cannot be written as a proper nontrivial power.
Lemma 2.2.
If $x,y,m\geq 2$ are integers and $x$ is a prime power, then the equation $x^{2}+x+1=y^{m}$ has no solution.
Proof.
The equation in the lemma can be rewritten as $(x^{3}-1)/(x-1)=y^{m}$. This is a special case of the Nagell-Ljunggren equation $(x^{n}-1)/(x-1)=y^{m}.$ Nagell and Ljunggren show that (see [3]) if $n=3$, then the only solution of the equation with $x,y,m\geq 2$ is $(x,y)=(18,7)$. However, $x=18$ is not a prime power. ∎
The next result is crucial to our proof of Step $3$ Unfortunately, we are unable to find a simple proof of this fact. The proof given below is an easy consequence of a deep result due to R. Guralnick and P.H. Tiep [8] concerning the non-coprime $k(GV)$-problem. For a finite group $G,$ we let $b(G)$ be the largest degree of $G.$ It follows from character theory that $|G|=\sum_{\chi\in{\mathrm{Irr}}(G)}\chi(1)^{2}$ and $k(G)=|{\mathrm{Irr}}(G)|,$ where $k(G)$ is the number of conjugacy classes of $G.$ Thus $|G|\leq k(G)b(G)^{2}$ and $b(G)^{2}\leq|G|.$
Lemma 2.3.
Let $L$ be a finite group and let $M$ be a minimal normal elementary abelian subgroup of $L.$ Assume that $L/M$ is isomorphic to a simple exceptional group of Lie type of Lie rank at least $4$ and that $C_{L}(M)=M.$ Then $b(L)>b(L/M).$
Proof.
Let $H\cong L/M$ be a finite simple exceptional group of Lie type of Lie rank at least $4$ By way of contradiction, assume that $b(L)\leq b(L/M)=b(H).$ Form the semidirect product $HM$ with $H$ acting on $M$ as inside $L.$ It follows from [8, Lemma 2.4] that $k(L)\leq k(HM).$ We now have that $|L|=\sum_{\chi\in{\mathrm{Irr}}(L)}\chi(1)^{2}\leq k(L)b(L)^{2}\leq k(HM)b(H)^{2}.$ As $H$ is a finite simple exceptional group of Lie type of Lie rank at least $4,$ and $H$ acts irreducibly on $M,$ by applying [8, Theorems 1.4, 1.5], we have $k(HM)\leq|M|/2.$ Furthermore $|L|=|H|\cdot|M|$ and so the above inequality becomes $|L|=|H|\cdot|M|\leq|M|b(H)^{2}/2.$ After simplifying, we obtain that $|H|\leq b(H)^{2}/2<b(H)^{2},$ which is impossible. Therefore, $b(L)>b(L/M)$ as wanted. ∎
To verify Step 3 for classical groups, we will need the following result.
Problem 2.4.
Let $G$ be a finite perfect group and let $M$ be a minimal normal elementary abelian subgroup of $G$ such that $G/M$ is a finite simple classical group. If $C_{G}(M)=M$, then $G$ has an irreducible character $\chi$ such that $\chi(1)$ divides no character degree of $G/M.$
It is even possible that $b(G)>b(G/M)$ in the above situation. To finish Step $3$, we still need to show that if $L$ is a quasisimple group with $L/Z(L)$ a finite simple classical group, then some character degree of $L$ divides no degree of $L/Z(L)$.
Finally, we need the following result due to Zsigmondy.
Lemma 2.5.
([15, Theorems  5.2.14, 5.2.15]). Let $q$ and $n$ be integers with $q\geq 2,$ and $n\geq 3.$ Suppose that $(q,n)\neq(2,6).$ Then $q^{n}-1$ has a prime divisor $\ell$ such that $\ell\equiv 1$ (mod $n$) and $\ell$ does not divide $q^{k}-1$ for any $k<n.$ Moreover if $\ell\mid q^{k}-1$ for some integer $k,$ then $n\mid k.$
Such a prime $\ell$ is called a primitive prime divisor (ppd for short) and is denoted by $\ell_{n}(q).$ When $n$ is odd and $(q,n)\neq(2,3)$ then there is a primitive prime divisor of $q^{2n}-1$ and we denote this primitive prime divisor by $l_{-n}(q).$ In terms of the cyclotomic polynomials, as $q^{n}-1=\prod_{k\mid n}\Phi_{k},$ we obtain that if $\ell$ is a primitive prime divisor of $q^{n}-1,$ then $\ell\mid\Phi_{n}$ but $\ell\nmid\Phi_{k}$ for any $k<n.$ In this situation, we also say that $\ell$ is a primitive prime divisor of $\Phi_{n}.$
Table 1. Character degrees of sporadic simple groups and the Tits group
Group Character Degree Group Character Degree
$\textrm{M}_{11}$ $\chi_{8}$ $2^{2}\cdot 11$ O’N $\chi_{2}$ $2^{6}\cdot 3^{2}\cdot 19$
$\chi_{9}$ $3^{2}\cdot 5$ $\chi_{19}$ $7^{3}\cdot 11\cdot 31$
$\textrm{M}_{12}$ $\chi_{7}$ $2\cdot 3^{3}$ $\textrm{Co}_{3}$ $\chi_{3}$ $11\cdot 23$
$\chi_{8}$ $5\cdot 11$ $\chi_{6}$ $2^{7}\cdot 7$
$\textrm{J}_{1}$ $\chi_{5}$ $2^{2}\cdot 19$ $\textrm{Co}_{2}$ $\chi_{3}$ $11\cdot 23$
$\chi_{6}$ $7\cdot 11$ $\chi_{22}$ $3^{6}\cdot 5^{3}$
$\textrm{M}_{22}$ $\chi_{2}$ $3\cdot 7$ $\textrm{Fi}_{22}$ $\chi_{56}$ $2^{17}\cdot 11$
$\chi_{5}$ $5\cdot 11$ $\chi_{57}$ $3^{9}\cdot 7\cdot 13$
$\textrm{J}_{2}$ $\chi_{6}$ $2^{2}\cdot 3^{2}$ HN $\chi_{10}$ $3^{4}\cdot 11\cdot 19$
$\chi_{13}$ $5^{2}\cdot 7$ $\chi_{45}$ $2^{10}\cdot 5^{5}$
$\textrm{M}_{23}$ $\chi_{3}$ $3^{2}\cdot 5$ Ly $\chi_{7}$ $2^{8}\cdot 7\cdot 67$
$\chi_{9}$ $11\cdot 23$ $\chi_{50}$ $3\cdot 5^{6}\cdot 31\cdot 37$
HS $\chi_{2}$ $2\cdot 11$ Th $\chi_{2}$ $2^{3}\cdot 31$
$\chi_{7}$ $5^{2}\cdot 7$ $\chi_{7}$ $5^{3}\cdot 13\cdot 19$
$\textrm{J}_{3}$ $\chi_{6}$ $2^{2}\cdot 3^{4}$ $\textrm{Fi}_{23}$ $\chi_{4}$ $13\cdot 17\cdot 23$
$\chi_{13}$ $5\cdot 17\cdot 19$ $\chi_{94}$ $2^{18}\cdot 5^{2}\cdot 7\cdot 11$
$\textrm{M}_{24}$ $\chi_{7}$ $2^{2}\cdot 3^{2}\cdot 7$ $\textrm{Co}_{1}$ $\chi_{3}$ $13\cdot 23$
$\chi_{8}$ $11\cdot 23$ $\chi_{17}$ $2\cdot 5^{4}\cdot 7^{2}\cdot 11$
McL $\chi_{2}$ $2\cdot 11$ $\textrm{J}_{4}$ $\chi_{2}$ $31\cdot 43$
$\chi_{14}$ $3^{6}\cdot 7$ $\chi_{11}$ $2^{3}\cdot 3^{2}\cdot 23\cdot 29\cdot 37$
He $\chi_{9}$ $3\cdot 5^{2}\cdot 17$ $\textrm{Fi}_{24}^{\prime}$ $\chi_{2}$ $13\cdot 23\cdot 29$
$\chi_{15}$ $2^{7}\cdot 7^{2}$ $\chi_{6}$ $5^{2}\cdot 7^{3}\cdot 11\cdot 17$
Ru $\chi_{5}$ $3^{3}\cdot 29$ B $\chi_{2}$ $3\cdot 31\cdot 47$
$\chi_{20}$ $2^{2}\cdot 5^{3}\cdot 7\cdot 13$ $\chi_{119}$ $2^{39}\cdot 11\cdot 19\cdot 23$
Suz $\chi_{2}$ $11\cdot 13$ M $\chi_{2}$ $47\cdot 59\cdot 71$
$\chi_{43}$ $2^{10}\cdot 3^{5}$ $\chi_{16}$ $5^{9}\cdot 7^{6}\cdot 11^{2}\cdot 17\cdot 19$
${}^{2}\textrm{F}_{4}(2)^{\prime}$ $\chi_{8}$ $5^{2}\cdot 13$
$\chi_{20}$ $2^{6}\cdot 3^{3}$
3. Character degrees of simple exceptional groups of Lie type
In this section, we collect all the results concerning the character degrees of exceptional groups of Lie type that we need for the proof of Theorem Theorem. We first consider the following set up. Let $\mathcal{G}$ be a simple simply connected algebraic group defined over the algebraic closure of a finite field ${\mathbb{F}}_{p}$ of characteristic $p$ and let $F:\mathcal{G}\longrightarrow\mathcal{G}$ be a Frobenius morphism such that the fixed point group $L=\mathcal{G}^{F}$ is finite and $H\cong L/Z(L)$ is a finite simple groups. Let $\mathcal{G}^{*}$ be a group dual to $\mathcal{G}$ with a dual Frobenius map $F^{*}$ and let $H^{*}=({\mathcal{G}^{*}})^{F^{*}}.$ We also use the notation $H_{sc}$ for $L$ and $H_{ad}$ for $H^{*}.$
According to Lusztig, there is a natural partition
${\mathrm{Irr}}(L)=\bigcup_{\begin{subarray}{c}s\in H^{*}/\sim\\ s\>\text{semisimple}\end{subarray}}\mathcal{E}(L,s)$
of ${\mathrm{Irr}}(L),$ where the union runs over $H^{*}$-conjugacy classes of semisimple elements. The irreducible characters of $L$ in $\mathcal{E}(L,1)$ are call unipotent characters of $L.$ The unipotent character degrees of simple exceptional groups of Lie type are given explicitly in [4, 13.9]. Moreover for each $s,$ there is a bijection map
$\psi_{s}:\mathcal{E}(L,s)\longrightarrow\mathcal{E}({\mathbf{C}}_{H^{*}}(s),1)$
such that for any $\chi\in\mathcal{E}(L,s),$ the degree of $\chi$ is given by
$\chi(1)=|H^{*}:{\mathbf{C}}_{H^{*}}(s)|\psi_{s}(\chi)(1)$
(see for example [19]). The semisimple character corresponding to the conjugacy class with representative $s$ is the character $\chi_{s}$ such that $\psi_{s}(\chi_{s})$ is the principal character of ${\mathbf{C}}_{H^{*}}(s).$ In this case $\chi_{s}(1)=|H^{*}:{\mathbf{C}}_{H^{*}}(s)|.$
By results of Lusztig, we know that all the unipotent characters of $L$ are trivial on the center of $L,$ therefore they are exactly all the unipotent characters of $H\cong L/Z(L).$ Moreover, by result of G. Malle in [17], all unipotent characters of $H$ extend to $\mathop{\rm Aut}\nolimits(H),$ apart from some exceptions (see [17, Theorem 2.5]). This fact will be used frequently in the proofs of Steps $2$ and $3$ Some results on semisimple characters of $L$ can be found in [6, Lemma 2.5]. In Table 2, we list the $p$-parts of the degrees of some unipotent characters of finite simple groups of Lie type in characteristic $p$.
Table 2. Some unipotent characters of finite simple groups of Lie type
$S=S(p^{b})$ Symbol $p$-part of degree
$\textrm{L}_{n}^{\epsilon}(p^{b}),n\geq 3$ $(1^{n-2},2)$ $p^{b(n-1)(n-2)/2}$
[…]
$\textrm{S}_{2n}(p^{b}),p>2$ $p^{b(n-1)^{2}}$
[…]
${}^{3}\textrm{D}_{4}(p^{b})$ $\phi_{1,3}^{\prime\prime}$ $p^{7b}$
$\textrm{F}_{4}(p^{b})$ $\phi_{9,10}$ $p^{10b}$
${}^{2}\textrm{F}_{4}(q^{2})$ ${}^{2}B_{2}[a],\epsilon$ $\frac{1}{\sqrt{2}}q^{13}$
$\textrm{E}_{6}(p^{b})$ $\phi_{6,25}$ $p^{25b}$
${}^{2}\textrm{E}_{6}(p^{b})$ $\phi_{2,16}^{\prime\prime}$ $p^{25b}$
$\textrm{E}_{7}(p^{b})$ $\phi_{7,46}$ $p^{46b}$
$\textrm{E}_{8}(p^{b})$ $\phi_{8,91}$ $p^{91b}$
All the results in this section are based on the data given in [16]. From this, we can extract all the information about the character degrees of both $L$ and $H^{*},$ the simply connected and adjoint versions of the simple exceptional group of Lie type $H.$ Note that if $d:=|Z(L)|=|H^{*}:H|$ is trivial, then all these versions of $H$ are isomorphic and we will obtain the character degrees of the simple group $H.$ Otherwise, we only know that ${\mathrm{cd}}(H)\subseteq{\mathrm{cd}}(L)$ and every character degree of $H$ must divide some degree of $H^{*}.$
We are now ready to obtain the results on the character degree sets of finite simple exceptional groups of Lie type. From now on we assume that $q=p^{f}$ is a power of a prime $p,$ where $f\geq 1$ and $H$ is isomorphic to one of the following simple groups:
$\textrm{F}_{4}(q),~{}{}^{2}\textrm{E}_{6}(q),~{}\textrm{E}_{6}(q),~{}\textrm{E}_{7}(q),~{}\textrm{E}_{8}(q),$ (1)
We first begin with the family $\textrm{F}_{4}(q).$ As the case when $H\cong\textrm{F}_{4}(2)$ has been handled in [29], we always assume that $q>3$ when $H\cong\textrm{F}_{4}(q).$
Lemma 3.1.
Let $H\cong{\rm F}_{4}(q),$ where $q\geq 3,$ and let $x\in{\mathrm{cd}}(H)$ with $x\not\in\{1,q^{24}\}$. Let $\ell_{i},i=1,2,\cdots,4,$ be primitive prime divisors of $\Phi_{12},\Phi_{8},\Phi_{6}$ and $\Phi_{3},$ respectively. Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2}.$
  2. $(ii)$
    If $(\ell_{1},x)=1,$ then $p\mid x$ or $x=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}.$
  3. $(iii)$
    If $(\ell_{2},x)=1,$ then $p\mid x$ or $x$ is $\Phi_{3}\Phi_{6}\Phi_{12}~{}(q~{}odd)$ or $\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{12}.$
  4. $(iv)$
    If $(\ell_{3}\ell_{4},x)=1,$ then $x=\frac{1}{2}q\Phi_{4}\Phi_{8}\Phi_{12},$ $q^{3}\Phi_{4}^{2}\Phi_{8}\Phi_{12},$ $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8},$ $q^{9}\Phi_{4}^{2}\Phi_{8}\Phi_{12},$ or $\frac{1}{2}q^{13}\Phi_{4}\Phi_{8}\Phi_{12}.$
  5. $(v)$
    There exist $\chi_{i}\in{\mathrm{Irr}}(H),i=1,2,$ with degrees $\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}$ and $\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{12},$ respectively, such that no proper multiple of each degree is a degree of $H.$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{16}$ when $q$ is odd and $\frac{1}{2}q^{13}$ when $q$ is even.
  7. $(vii)$
    The smallest nontrivial character degree of $H$ is $\Phi_{3}\Phi_{6}\Phi_{12}$ when $q$ is odd and $\frac{1}{2}q\Phi_{1}^{2}\Phi_{3}^{2}\Phi_{8}$ when $q$ is even.
  8. $(viii)$
    Assume $x,y\in{\mathrm{cd}}(H)-\{1,q^{24}\}.$ If $p\neq 3,$ then $(x,y)>1.$ If $p=3$ and $(x,y)=1,$ then $\{x,y\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$
  9. $(ix)$
    $H$ has no consecutive degrees.
Proof.
Statements $(i)-(vii)$ can be checked directly given the character degree set of $H$ in [16]. We only need to verify statements $(viii)$ and $(ix).$ For $(viii),$ assume that $x,y\in{\mathrm{cd}}(H)-\{1,q^{24}\}$ with $(x,y)=1.$ By $(i),$ we have that either $(\ell_{1}\ell_{2},x)>1$ or $(\ell_{1}\ell_{2},y)>1.$ Without loss of generality, we assume that $(\ell_{1}\ell_{2},x)>1.$
Assume first that $\ell_{1}\mid x.$ Then $(\ell_{1},y)=1.$ By $(ii),$ either $y=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}$ or $p\mid y.$ If the first case holds, then $\Phi_{3}\Phi_{4}\Phi_{6}\Phi_{8}\mid y$ and so as $(x,y)=1$ we deduce that $(\Phi_{3}\Phi_{4}\Phi_{6}\Phi_{8},x)=1,$ which is impossible by $(iv).$ Thus $p\mid y$ and hence $p\nmid x.$ If $(\ell_{2},y)=1,$ then $y$ is coprime to both $\ell_{1}\ell_{2}$ so $y=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2}$ by $(i).$ But then $\Phi_{3}\Phi_{6}\mid y$ and so $(\Phi_{3}\Phi_{6},x)=1.$ Moreover as $p\nmid x,$ we deduce that $(p\Phi_{3}\Phi_{6},x)=1,$ which is impossible by $(iv).$ Thus $\ell_{2}\mid y$ and hence $(\ell_{2},x)=1,$ so $(p\ell_{2},x)=1.$ It follows from $(iii)$ that $x=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{12}$ or $q$ is odd and $x=\Phi_{3}\Phi_{6}\Phi_{12}.$ If the former case holds then $\Phi_{3}\Phi_{4}\Phi_{6}\mid x$ and so $(\Phi_{3}\Phi_{4}\Phi_{6},y)=1,$ which is impossible by $(iv).$ Therefore $q$ is odd and $x=\Phi_{3}\Phi_{6}\Phi_{12}.$ Using $(iv)$ again, the only possibility for $y$ is $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}.$ If $p=3,$ then we are done. Now assume that $p\neq 3.$ We will show that $(\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8})>1.$ As $q$ is coprime to $3,$ we deduce that $\Phi_{1}^{4}\Phi_{2}^{4}=(q^{2}-1)^{4}$ is divisible by $3^{4}$ and thus $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}$ is divisible by $3$ Moreover $\Phi_{3}\Phi_{6}\Phi_{12}=q^{8}+q^{4}+1=(q^{8}-1)+(q^{4}-1)+3$ is also divisible by $3,$ therefore the greatest common divisor of $\Phi_{3}\Phi_{6}\Phi_{12}$ and $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}$ is divisible by $3$ so it is nontrivial.
Assume that $\ell_{2}\mid x.$ Then $(\ell_{2},y)=1.$ If $(\ell_{1},y)=1,$ then $y=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2}$ by $(i).$ Hence $p\Phi_{3}\Phi_{6}\mid y$ so $(p\Phi_{3}\Phi_{6},x)=1,$ which is impossible by $(iv).$ Thus $\ell_{1}\mid y$ and so $(\ell_{1},x)=1.$ By $(ii),$ we obtain that $x=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}$ or $p\mid x.$ If the former case holds, then $\Phi_{3}\Phi_{4}\Phi_{6}\Phi_{8}\mid x$ and so $(\Phi_{3}\Phi_{4}\Phi_{6}\Phi_{8},y)=1,$ which is impossible by $(iv).$ Thus $p\mid x$ and hence $p\nmid y.$ Thus $(p\ell_{2},y)=1.$ It follows from $(iii)$ that $y=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{12}$ or $q$ is odd and $y=\Phi_{3}\Phi_{6}\Phi_{12}.$ Using the same argument as in the previous case, we obtain the required result.
We next show that $H$ has no consecutive degrees. Assume that $x=y+1.$ By $(vii),$ we can assume that $x>y>1.$ Since $q^{24}\pm 1$ do not divide the order of $H,$ we deduce that $H$ has no such degrees. So, we also assume that $x,y\in{\mathrm{cd}}(H)-\{1,q^{24}\}.$ Moreover $(x,y)=1$ and so by $(viii),$ we must have that $p=3$ and $\{x,y\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$ Observe that $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}>q^{14}>q^{9}>\Phi_{3}\Phi_{6}\Phi_{12}$ as $q\geq 3,$ therefore $x=\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}~{}\mbox{and}~{}y=\Phi_{3}\Phi_{6}\Phi_{12}.$ But then $y+1<q^{9}+1<q^{14}<x,$ a contradiction. The proof is now complete. ∎
For the remaining lemmas, the proofs are basically identical to that of Lemma 3.1, so we will skip all the detail.
Lemma 3.2.
Let $H$ be the simple exceptional group ${}^{2}{\rm E}_{6}(q)$ and let $\ell_{i},i=1,2,3,4$ be primitive prime divisors of $\Phi_{18},\Phi_{12},\Phi_{8}$ and $\Phi_{10},$ respectively. Let $x\neq q^{36}$ be a nontrivial character degree of $L\cong{}^{2}{\rm E}_{6}(q)_{sc}.$ Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbols ${}^{2}E_{6}[\theta^{i}],i=1,2$ with degree $\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}.$
  2. $(ii)$
    If $(\ell_{2}\ell_{3},x)=1,$ then $x$ is the degree of the unipotent characters labeled by the symbols $\phi_{8,3}^{\prime}$ or $\phi_{8,9}^{\prime\prime}$ with degree $\frac{1}{2}q^{3}\Phi_{2}^{4}\Phi_{6}^{2}\Phi_{10}\Phi_{18}~{}\mbox{or~{}}~{}\frac{1}{2}q^{15}\Phi_{2}^{4}\Phi_{6}^{2}\Phi_{10}\Phi_{18}.$
  3. $(iii)$
    We have $(\ell_{1}\ell_{3},x)>1$ and $(p\ell_{1}\ell_{4},x)>1.$
  4. $(iv)$
    If $q=2$ then $(p\ell_{3}\ell_{4},x)>1,$ and if $q>2$ and $(p\ell_{3}\ell_{4},x)=1,$ then $x=\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}.$
  5. $(v)$
    $H$ possesses two semisimple irreducible characters $\chi_{i},i=1,2$ corresponding to the maximal tori with orders $\Phi_{18}$ and $\Phi_{12}\Phi_{6},$ with degrees
    $\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{3}\Phi_{8}\Phi_{10}\Phi_{12},\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}\Phi_{10}\Phi_{18},$
    respectively, such that no proper multiple of each degree is a degree of $H.$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{25}.$
  7. $(vii)$
    The unipotent character labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $q\Phi_{8}\Phi_{18}$ is the smallest nontrivial character degree of $H.$
  8. $(viii)$
    Assume $x,y\in{\mathrm{cd}}(L)-\{1,q^{36}\}.$ If $p\neq 3,$ then $(x,y)>1.$ If $p=3$ and $(x,y)=1,$ then $\{x,y\}=\{\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10},~{}\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}\}$.
  9. $(ix)$
    $L$ has no consecutive degrees.
  10. $(x)$
    Suppose that $13\mid q+$ 1. There exists $\chi\in{\mathrm{Irr}}(L)$ of degree $\frac{1}{3}q^{9}\Phi_{1}^{3}\Phi_{2}^{4}\Phi_{3}\Phi_{4}^{2}\Phi_{6}\Phi_{8}\Phi_{10}\Phi_{12}$ with multiplicity $6$ such that $\chi(1)$ divides no degree of $H.$ Similarly $H^{*}\cong{}^{2}{\rm E}_{6}(q)_{ad}$ has an irreducible character $\psi$ with $\psi(1)=q^{9}\Phi_{1}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\Phi_{12}\Phi_{18}$ with multiplicity $1$ such that $\psi(1)$ is not a degree of $H.$
Lemma 3.3.
Let $H$ be the simple exceptional group ${\rm E}_{6}(q)$ and let $\ell_{i},i=1,2,3,4$ be primitive prime divisors of $\Phi_{12},\Phi_{9},\Phi_{8}$ and $\Phi_{5},$ respectively. Let $x\neq q^{36}$ be a nontrivial character degree of $L.$ Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbols $E_{6}[\theta^{i}],i=1,2,$ with degree $\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{5}\Phi_{8}.$
  2. $(ii)$
    We have $(x,\ell_{2}\ell_{3})>1.$
  3. $(iii)$
    If $(x,\ell_{1}\ell_{3})=1,$ then $x$ is the degree of unipotent characters labeled by the symbols $(D_{4},1)$ or $(D_{4},\epsilon)$ with degrees $\frac{1}{2}q^{3}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}\Phi_{9}~{}\mbox{or~{}}\frac{1}{2}q^{15}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}\Phi_{9}.$
  4. $(iv)$
    We have $(x,p\ell_{2}\ell_{4})>1,$ and if $(x,p\ell_{3}\ell_{4})=1$ then $x=\Phi_{3}^{2}\Phi_{6}\Phi_{9}\Phi_{12}.$
  5. $(v)$
    $H$ possesses two irreducible characters $\chi_{i},i=1,2$ with degrees $\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{5}\Phi_{8}\Phi_{9}$ and $\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{3}^{3}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{5}\Phi_{8}\Phi_{12},$ respectively such that no proper multiples of each degree is a degree of $H.$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{25}.$
  7. $(vii)$
    The unipotent character labeled by the symbol $\phi_{6,1}$ with degree $q\Phi_{8}\Phi_{9}$ is the smallest nontrivial character degree of $H.$
  8. $(viii)$
    Assume $x,y\in{\mathrm{cd}}(L)-\{1,q^{36}\}.$ If $p\neq 3,$ then $(x,y)>1.$ If $p=3$ and $(x,y)=1,$ then $\{x,y\}=\{\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{5}\Phi_{8},~{}\Phi_{3}^{2}\Phi_{6}\Phi_{9}\Phi_{12}\}.$
  9. $(ix)$
    $L$ has no consecutive degrees.
  10. $(x)$
    Assume $13\mid q-$ 1 . There exists $\chi\in{\mathrm{Irr}}(L)$ of degree $\frac{1}{3}q^{9}\Phi_{1}^{4}\Phi_{2}^{3}\Phi_{3}\Phi_{4}^{2}\Phi_{5}\Phi_{6}\Phi_{8}\Phi_{12}$ with multiplicity $6$ such that $\chi(1)$ divides no degree of $H.$ Moreover ${\rm E}_{6}(q)_{ad}$ has an irreducible character $\psi$ of degree $q^{9}\Phi_{2}\Phi_{4}^{2}\Phi_{5}\Phi_{6}^{2}\Phi_{8}\Phi_{9}\Phi_{12}$ with multiplicity $1$ such that $\psi(1)$ is not a degree of $H.$
Lemma 3.4.
Let $H$ be the simple exceptional group ${\rm E}_{7}(q)$ and let $\ell_{i},i=1,2,\cdots,5$ be primitive prime divisors of $\Phi_{18},\Phi_{14},$ $\Phi_{12},\Phi_{9},$ and $\Phi_{7},$ respectively. Let $x$ be a nontrivial character degree of $L$ with $x\neq q^{63}.$ Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbol $E_{7}[\xi]$ or $E_{7}[-\xi]$ with degree $\frac{1}{2}q^{11}\Phi_{1}^{7}\Phi_{3}^{3}\Phi_{4}^{2}\Phi_{5}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{12}.$
  2. $(ii)$
    If $(\ell_{2}\ell_{3},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbols $(D_{4},\epsilon_{1})$ or $(D_{4},\epsilon_{2})$ with degree
    $\frac{1}{2}q^{4}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}\Phi_{7}\Phi_{9}\Phi_{10}\Phi_{18}\text{ or }\frac{1}{2}q^{25}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}\Phi_{7}\Phi_{9}\Phi_{10}\Phi_{18}.$
  3. $(iii)$
    If $(\ell_{1}\ell_{3},x)=1,$ then $\ell_{4}\ell_{5}\mid x$ or $x$ is the degree of unipotent characters labeled by $(E_{6}[\theta^{i}],1)$ or $(E_{6}[\theta^{i}],\epsilon),$ where $i=1,2,$ with degrees
    $\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{5}\Phi_{7}\Phi_{8}\Phi_{10}\Phi_{14}\text{ or }\frac{1}{3}q^{16}\Phi_{1}^{6}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{5}\Phi_{7}\Phi_{8}\Phi_{10}\Phi_{14}.$
  4. $(iv)$
    We have $(p\ell_{2}\ell_{5},x)>1,$ and if $(\ell_{4}\ell_{5},x)=1,$ then $x$ is the degree of the unipotent characters labeled by $\phi_{512,11}$ or $\phi_{512,12}$ with degree $\frac{1}{2}q^{11}\Phi_{2}^{7}\Phi_{4}^{2}\Phi_{6}^{3}\Phi_{8}\Phi_{10}\Phi_{12}\Phi_{14}\Phi_{18}.$
  5. $(v)$
    $H$ possesses two semisimple irreducible characters $\chi_{i},i=1,2$ corresponding to the maximal tori with orders $\Phi_{18}\Phi_{2}$ and $\Phi_{14}\Phi_{2},$ respectively, and having degrees
    $\Phi_{1}^{7}\Phi_{2}^{6}\Phi_{3}^{3}\Phi_{4}^{2}\Phi_{5}\Phi_{6}^{3}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{10}\Phi_{12}\Phi_{14}\text{ and }\Phi_{1}^{7}\Phi_{2}^{6}\Phi_{3}^{3}\Phi_{4}^{2}\Phi_{5}\Phi_{6}^{3}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{10}\Phi_{12}\Phi_{18},$
    respectively such that no proper multiples of these degrees is in ${\mathrm{cd}}(L).$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{46}.$
  7. $(vii)$
    The unipotent character labeled by the symbol $\phi_{7,1}$ with degree $q\Phi_{7}\Phi_{12}\Phi_{14}$ is the smallest nontrivial character degree of $H.$
  8. $(viii)$
    If $x,y\in{\mathrm{cd}}(L)-\{1,q^{63}\},$ then $(x,y)>1.$
  9. $(ix)$
    $L$ has no consecutive degrees.
  10. $(x)$
    If $q$ is odd and $q\equiv\epsilon$ (mod $4$), where $q\in\{\pm 1\},$ then $L$ has an irreducible character $\chi$ of degree $q^{14}\Phi_{1}^{3}\Phi_{2}^{2}\Phi_{3}^{2}\Phi_{5}\Phi_{6}^{2}\Phi_{7}\Phi_{9}\Phi_{10}\Phi_{12}\Phi_{14}\Phi_{18}$ with multiplicity ${(q-\epsilon)}/{4}$ such that $\chi(1)$ divides no degree of $H.$ Moreover $E_{7}(q)_{ad}$ has a nontrivial degree $\psi(1)$ which is not a degree of $H.$ In particular, if $q\equiv 1$ (mod 4), then we can choose
    $\psi(1)=q^{28}\Phi_{2}^{3}\Phi_{3}\Phi_{6}^{2}\Phi_{9}\Phi_{10}\Phi_{12}\Phi_{14}\Phi_{18}$
    and if $q\equiv-1$ (mod 4), then $\psi(1)=q^{28}\Phi_{1}^{3}\Phi_{3}^{2}\Phi_{5}\Phi_{6}\Phi_{7}\Phi_{9}\Phi_{12}\Phi_{18}.$
Lemma 3.5.
Let $H$ be the simple exceptional group ${\rm E}_{8}(q)$ and let $\ell_{i},i=1,2,\cdots,5,$ be primitive prime divisors of $\Phi_{30},\Phi_{24},\Phi_{20},\Phi_{15}$ and $\Phi_{14},$ respectively. Let $x$ be a nontrivial character degree of $H$ with $x\neq q^{120}.$ Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbol $E_{8}[-\theta]$ or $E_{8}[-\theta^{2}]$ with degree $\frac{1}{6}q^{16}\Phi_{1}^{8}\Phi_{2}^{6}\Phi_{3}^{2}\Phi_{4}^{4}\Phi_{5}^{2}\Phi_{7}\Phi_{8}^{2}\Phi_{9}\Phi_{10}^{2}\Phi_{12}\Phi_{14}\Phi_{15}\Phi_{20}.$
  2. $(ii)$
    If $(\ell_{2}\ell_{3},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbol $E_{8}[i]$ or $E_{8}[-i]$ with degree $\frac{1}{4}q^{16}\Phi_{1}^{8}\Phi_{2}^{8}\Phi_{3}^{4}\Phi_{5}^{2}\Phi_{6}^{4}\Phi_{7}\Phi_{9}\Phi_{10}^{2}\Phi_{14}\Phi_{15}\Phi_{18}\Phi_{30}.$
  3. $(iii)$
    If $(\ell_{1}\ell_{3},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbol $E_{8}[\zeta^{k}],k=1,2,\cdots,4$ with degree $\frac{1}{5}q^{16}\Phi_{1}^{8}\Phi_{2}^{8}\Phi_{3}^{4}\Phi_{4}^{4}\Phi_{6}^{4}\Phi_{7}\Phi_{8}^{2}\Phi_{9}\Phi_{12}^{2}\Phi_{14}\Phi_{18}\Phi_{24}$ or $x$ is the degree of unipotent characters labeled by the symbol $(D_{4},\phi_{1,0})$ or $(D_{4},\phi_{1,24})$ with degrees $\frac{1}{2}q^{3}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}^{2}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{14}\Phi_{15}\Phi_{24},~{}\frac{1}{2}q^{63}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}^{2}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{14}\Phi_{15}\Phi_{24},$ respectively.
  4. $(iv)$
    We have $(p\ell_{4}\ell_{5},x)>1,$ and $(p\ell_{2}\ell_{5},x)>1.$
  5. $(v)$
    $H$ has irreducible characters $\chi_{i},i=1,2,3$ with degrees $|H|/\Phi_{m_{i}},$ where $m_{i}$ is $30,24$ and $20,$ respectively. Moreover no proper multiples of $\chi_{i}(1),i=1,2,3$ is a degree of $H.$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{91}.$
  7. $(vii)$
    The unipotent character with degree $q\Phi_{4}^{2}\Phi_{8}\Phi_{12}\Phi_{20}\Phi_{24}$ labeled by the symbol $\phi_{8,1}$ is the smallest nontrivial character degree of $H.$
  8. $(viii)$
    If $x,y\in{\mathrm{cd}}(H)-\{1,q^{120}\},$ then $(x,y)>1.$
  9. $(ix)$
    $H$ has no consecutive degrees.
4. Verifying Step 1: $G^{\prime}=G^{\prime\prime}$
In this section, we will show that $G^{\prime}=G^{\prime\prime}.$ By way of contradiction, suppose that $G^{\prime}\neq G^{\prime\prime}.$ Then there exists a normal subgroup $N\unlhd G$ of $G$ such that $G/N$ is solvable minimal with respect to being nonabelian. By [25, Lemma 2.3], $G/N$ is an $r$-group for some prime $r$ or $G/N$ is a Frobenius group.
Recall that if $\chi\in{\mathrm{Irr}}(G)$, then we say that $\chi$ is isolated in $G$ if $\chi(1)$ is divisible by no proper nontrivial character degree of $G,$ and no proper multiple of $\chi(1)$ is a character degree of $G.$ In this situation, we also say that $\chi(1)$ is an isolated degree of $G.$
Case $1$ $G/N$ is an $r$-group. By [18, Theorem $1.1$], the only nontrivial prime power degree of $H$ is the degree of the Steinberg character of $H,$ which is $|H|_{p}.$ Since $G/N$ is a nonabelian $r$-group, it possesses an irreducible character $\tau$ of degree $r^{b}$ so $\tau(1)=r^{b}=|H|_{p}$ and hence $r=p.$ By Thompson’s Theorem ([13, Corollary 12.2]), $H$ has a nontrivial irreducible character $\psi$ with $p\nmid\psi(1),$ so by [9, Lemma $2$], we have that $\psi_{N}\in{\mathrm{Irr}}(N)$ and hence by Gallagher’s Theorem, $\tau\psi$ is an irreducible character of $G$ of degree $\tau(1)\psi(1)=|H|_{p}\psi(1).$ But this contradicts [25, Lemma 2.4].
Case $2$ $G/N$ is a Frobenius group with kernel $F/N,$ $|F/N|=r^{a},$ $1<m=|G:F|\in{\mathrm{cd}}(G)$ and $r^{a}\equiv 1(\mbox{mod $m$}).$ By [25, Lemma 2.4], we have that ${\mathrm{St}}_{H}(1)=|H|_{p}$ is an isolated degree of $H$ so by [25, Lemma 2.3(b)(2)], either $m=|H|_{p}$ or $r=p.$ We consider each family of simple groups separately.
Subcase $H\cong{\rm F}_{4}(q).$ Recall that $q>2$ in this case. Assume first that $m=|H|_{p}.$ As $r\nmid m,$ we have that $r\neq p.$ By [4, $13.9$], $H$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $\varphi(1)=\frac{1}{2}q\Phi_{4}\Phi_{8}\Phi_{12}.$ As no proper multiple of $m$ is a character of $G$ and $\varphi(1)\nmid m,$ we deduce from [25, Lemma 2.3(b)(1)] that $r^{a}\mid\varphi(1)^{2}.$ As $r\neq p,$ we obtain that $r^{a}\mid\varphi(1)_{p^{\prime}}^{2}=\frac{1}{4}\Phi_{4}^{2}\Phi_{8}^{2}\Phi_{12}^{2}.$ We have $\frac{1}{2}\Phi_{4}\Phi_{8}\Phi_{12}=\frac{1}{2}(q^{4}+1)(q^{6}+1)<q^{10}.$ Hence $r^{a}\leq\varphi(1)_{p^{\prime}}^{2}<q^{20}.$ Since $m\mid r^{a}-1,$ $m=q^{24}\leq r^{a}-1<q^{20}-1<q^{20},$ a contradiction.
Thus $r=p.$ Hence $m\in{\mathrm{cd}}(G)$ is coprime to $p$ and so $m\neq|H|_{p}.$ Let $\chi_{i}\in{\mathrm{Irr}}(H),i=1,2,$ be irreducible characters of $H$ obtained from Lemma 3.1$(v)$. For each $i,$ as no proper multiple of $\chi_{i}(1)$ is a degree of $H,$ [25, Lemma 2.3(b)(2)] implies that $m$ divides $\chi_{i}(1)$ and hence $m$ divide $\gcd(\chi_{1}(1),\chi_{2}(1)),$ which implies that $m$ is coprime to $\ell_{1}\ell_{2}.$ By Lemma 3.1$(i)$, $m=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2},$ which is impossible as $m$ is coprime to $p.$
Subcase $H\cong{}^{2}{\rm E}_{6}(q).$ Assume first that $m=|H|_{p}.$ As $r\nmid m,$ $r\neq p.$ By Lemma 3.2$(vii)$, $H$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $\varphi(1)=q\Phi_{8}\Phi_{18}.$ As no proper multiple of $m$ is a character of $G$ and $\varphi(1)\nmid m,$ we deduce from [25, Lemma 2.3(b)(1)] that $r^{a}\mid\varphi(1)^{2}.$ As $r\neq p,$ we obtain that $r^{a}\mid\varphi(1)_{p^{\prime}}^{2}=\Phi_{8}^{2}\Phi_{18}^{2}.$ We have $\Phi_{8}\Phi_{18}=q^{10}-q^{7}+q^{6}+q^{4}-q^{3}+1<q^{15}.$ As $r^{a}\mid\varphi(1)_{p^{\prime}}^{2},$ we deduce that $r^{a}\leq\varphi(1)_{p^{\prime}}^{2}<q^{30}.$ As $m\mid r^{a}-1,$ $m=q^{36}\leq r^{a}-1<q^{30}-1<q^{30},$ which is impossible.
Thus $r=p.$ Then $m\in{\mathrm{cd}}(G)-\{1,|H|_{p}\}$ is coprime to $p.$ By [25, Lemma 2.3(b)(2)] and Lemma 3.2$(v),$ $m$ divides both $\chi_{i}(1),i=1,2.$ As $\chi_{i}(1)$ is coprime to $\ell_{i}$ for $i=1,2,$ we deduce that $m$ is coprime to $\ell_{1}\ell_{2}$ so by Lemma 3.2$(i)$, $m=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}$, which is impossible as $p\nmid m.$
The argument for the remaining cases are similar.
5. Verifying Step 2: $G^{\prime}/M$ is isomorphic to $H$
Let $M\leq G^{\prime}$ be a normal subgroup of $G$ such that $G^{\prime}/M$ is a chief factor of $G.$ As $G^{\prime}$ is perfect, $G^{\prime}/M$ is nonabelian so $G^{\prime}/M\cong S^{k}$ for some finite nonabelian simple group $S$ and some integer $k\geq 1.$
5.1. Eliminating the alternating groups of degree at least $7$
Assume that $S=\mathrm{A}_{n},n\geq 7.$ Let $\theta_{i},i=1,2$ be irreducible characters of $S$ obtained from Lemma 2.1$(i)$ with $\theta_{1}(1)=n(n-3)/2$ and $\theta_{2}(1)=(n-1)(n-2)/2.$ We have that both $\theta_{i}$ extend to ${\mathop{\rm Aut}\nolimits}(\mathrm{A}_{n})\cong\mathrm{S}_{n}$ and $\theta_{2}(1)=\theta_{1}(1)+1.$ By [2, Lemma 5], we obtain $\theta_{i}^{k}\in{\mathrm{Irr}}(G^{\prime}/M)$ extends to $G/M,$ for $i=1,2.$ Hence $\theta_{i}(1)^{k}\in{\mathrm{cd}}(G)$ and $\theta_{i}(1)^{k}$ are coprime for $i=1,2.$ Note that $\mathrm{S}_{n}$ also has an irreducible character labelled by the partition $(n-1,1)$ which is irreducible upon restriction to $\mathrm{A}_{n}$. Thus $\mathrm{A}_{n}$ has an irreducible character $\theta$ of degree $n-1$ which is extendible to $\mathop{\rm Aut}\nolimits(\mathrm{A}_{n})\cong\mathrm{S}_{n}$. Hence $\theta(1)^{k}=(n-1)^{k}\in{\mathrm{cd}}(G)$.
We will need the following observation:
$\Phi_{3}\Phi_{6}\Phi_{12}=q^{8}+q^{4}+1<\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}=\frac{1}{3}q^{4}(q^{2}-1)^{2}(q^{4}-1)(q^{8}-1).$
If $q=3^{f},$ where $f\geq 1,$ then $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}$ is divisible by $q^{8}-1=3^{8f}-1,$ which is in turn divisible by $3^{8}-1=2^{5}\cdot 5\cdot 41,$ and so $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}$ is always divisible by $2^{5}\cdot 3\cdot 5\cdot 41$
Subcase $H\cong\textrm{F}_{4}(q),q>2.$ Assume first that $p\neq 3.$ By Lemma 3.1$(viii),$ one of the degrees $\theta_{i}(1)^{k},i=1,2,$ must be $q^{24}.$ However we have that $(n-1,n-2)=1$ and $(n,n-3)=(n,3)$ so $\theta_{i}(1)^{k}$ can never be a power of $p\neq 3.$ Thus $p=3.$ If $\theta_{j}^{k}(1)=q^{24}$ for some $j,$ then $j=1$ and hence $\theta_{1}^{k}(1)=(n(n-3)/2)^{k}=3^{24f},$ where $q=3^{f}.$ As $n\geq 7,$ we see that $n(n-3)/2$ is a prime power if and only if $n=9,$ so that $\theta_{1}^{k}(1)=3^{3k}=3^{24f}.$ Now $G$ has a degree $(n-1)^{k}=8^{k}$. By [18, Theorem $1.1$], the only nontrivial prime power degree of $G$ is $q^{24}=3^{24f},$ whence $3^{24f}=8^{k}$, a contradiction. Now Lemma 3.1$(viii)$ yields
$\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$
By Lemma 3.1$(ix)$, $G$ has no consecutive degrees so $k\geq 2.$ But then $\Phi_{3}\Phi_{6}\Phi_{12}$ is a proper nontrivial power, which is impossible by Lemma 2.2.
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ Assume first that $p\neq 3.$ By Lemma 3.2$(viii),$ one of the degrees $\theta_{i}(1)^{k},i=1,2,$ must be $q^{36}.$ However we have that $(n-1,n-2)=1$ and $(n,n-3)=(n,3)$ so $\theta_{i}(1)^{k}$ can never be a power of $p\neq 3.$ Thus $p=3.$ If $\theta_{j}^{k}(1)=q^{36}$ for some $j,$ then $j=1$ and hence $\theta_{1}^{k}(1)=(n(n-3)/2)^{k}=3^{36f},$ where $q=3^{f}.$ As $n\geq 7,$ we see that $n(n-3)/2$ is a prime power if and only if $n=9,$ so that $\theta_{1}^{k}(1)=3^{3k}=3^{36f}.$ Again, $G$ has a degree $(n-1)^{k}=8^{k}$. However, by [18, Theorem $1.1$], the only nontrivial prime power degree of $G$ is $q^{36}=3^{36f},$ we obtain a contradiction. Thus by Lemma 3.2$(viii)$, we have $\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18},\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\}.$ By Lemma 3.2$(ix)$, $G$ has no consecutive degrees so $k\geq 2.$ But then $\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}$ is a proper nontrivial power. As $q=3^{f},$ we deduce that $\Phi_{3}$ is coprime to $\Phi_{6}^{2}\Phi_{12}\Phi_{18}$ and so $\Phi_{3}$ must be a nontrivial proper power, which is impossible by Lemma 2.2.
The arguments for the remaining groups $\textrm{E}_{6}(q),\textrm{E}_{7}(q$ and $\textrm{E}_{8}(q)$ are similar.
5.2. Eliminating the sporadic simple groups and the Tits group.
Suppose that $S$ is a sporadic simple group or the Tits group. By Table 1, $S$ possesses two nontrivial irreducible characters $\theta_{i},i=1,2,$ such that $\theta_{i}(1)^{\prime}s$ are coprime and each $\theta_{i}$ extends to $\mathop{\rm Aut}\nolimits(S).$ By [2, Lemma 5], for each $i,$ $\theta_{i}^{k}$ extends to $G,$ so $\theta_{i}^{k}(1)\in{\mathrm{cd}}(G).$ Observe that for each $i,$ $\theta_{i}(1)$ is not prime power so $\theta_{i}^{k}(1)\neq|H|_{p}.$
Assume $H\cong{\rm F}_{4}(q)$. As $(\theta_{1}^{k}(1),\theta_{2}^{k}(1))=1,$ we deduce from Lemma 3.1$(viii)$ that $q=3^{f}$ and $\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$ In this situation, the larger degree among $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
Assume $H\cong{}^{2}{\rm E}_{6}(q)$. As $(\theta_{1}^{k}(1),\theta_{2}^{k}(1))=1,$ we deduce from Lemma 3.2$(viii)$ that $q=3^{f}$ and $\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18},\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\}.$ In this situation, one of the degrees $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
Assume $H\cong{\rm E}_{6}(q)$. As $(\theta_{1}^{k}(1),\theta_{2}^{k}(1))=1,$ we deduce from Lemma 3.3$(viii)$ that $q=3^{f}$ and $\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}^{2}\Phi_{6}\Phi_{9}\Phi_{12},\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{5}\Phi_{8}\}.$ In this situation, one of the degrees $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
Assume $H\cong{\rm E}_{7}(q)$ or ${\rm E}_{8}(q)$. Observe that both $\theta_{i}(1)$ are not prime powers so both $\theta_{i}^{k}(1)$ are pairwise coprime and different from $|H|_{p},$ which contradicts Lemma 3.4$(viii)$ and Lemma 3.5$(viii)$, respectively.
5.3. Reduction to finite simple groups of Lie type in characteristic $p$
In this subsection, we show that if $S$ is a finite simple group of Lie type in characteristic $r,$ with $S\neq{}^{2}{\rm F}_{4}(2)^{\prime},$ then $k=1$ and $r=p.$ Let $\theta$ be the Steinberg character of $S.$ Then $\theta(1)=|S|_{r}$ and $\theta$ extends to ${\mathop{\rm Aut}\nolimits}(S).$ By [2, Lemma 5], $\theta^{k}\in{\mathrm{Irr}}(G^{\prime}/M)$ extends to $G/M,$ hence $\theta(1)^{k}=|S|_{r}^{k}\in{\mathrm{cd}}(G).$ Since $|H|_{p}$ is the unique nontrivial prime power character degree of $G$ by [18, Theorem $1.1$], we have $\theta(1)^{k}=|H|_{p}.$ In particular, we have $r=p.$
We next claim that $k=1$. By way of contradiction, assume that $k\geq 2.$ Write $\theta(1)=q_{1}^{s}$ and $|H|_{p}=q^{a(H)}$, so $a(H)=24,36,36,63,120$, respectively (see [15, Table 5.1.B]). Then $q_{1}^{sk}=q^{a(H)}.$ Let $\psi=\tau\times\theta\times\cdots\times\theta\in{\mathrm{Irr}}(S^{k}),$ where $\tau\in{\mathrm{Irr}}(S)$ with $1<\tau(1)\neq\theta(1).$ Then $\psi(1)=\tau(1)q_{1}^{s(k-1)}\nmid q^{a(H)}$ and is nontrivial, so it must divide some character degree of $G,$ which is different from $q^{a(H)}.$ By part $(vi)$ of Lemmas 3.13.5, we have that $q_{1}^{s(k-1)}=q^{a(H)(k-1)/k}\leq q^{b(H)}$, where $b(H)=16,25,25,46,91$, respectively, and hence $a(H)(k-1)\leq kb(H),$ which implies that $k\leq k_{0}$ for some positive integer $k_{0}$. We can check that $k_{0}=3$ for all groups except for ${\rm E}_{8}(q)$, where $k_{0}=4$.
Let $C$ be a normal subgroup of $G$ such that $C/M=C_{G/M}(G^{\prime}/M).$ Then $G^{\prime}C/C\cong S^{k}$ is a unique minimal normal subgroup of $G/C$ so $G/C$ embeds into ${\mathop{\rm Aut}\nolimits}(S)\wr S_{k},$ where $S_{k}$ is the symmetric group of degree $k.$ Let $B={\mathop{\rm Aut}\nolimits}(S)^{k}\cap G/C.$ Then $|G/C:B|\mid k!.$ Let $\psi=1\times\theta\times\cdots\times\theta\in{\mathrm{Irr}}(G^{\prime}C/C).$ Then $\psi$ extends to $\psi_{0}\in{\mathrm{Irr}}(B)$ by using the representation theory of wreath products. Let $J=I_{G}(\psi_{0})$ and let $\mu\in{\mathrm{Irr}}(J|\psi_{0}).$ By [13, Theorem 6.11], $\chi=\mu^{G}\in{\mathrm{Irr}}(G).$ Then $\chi(1)=e\psi_{0}(1)=e\psi(1)=eq^{a(H)(k-1)/k},$ where $e\mid k!.$ Since $k\leq 4,$ we deduce that $e\mid 4!=24,$ so $1<\chi(1)<q^{a(H)}.$ In addition $\chi(1)$ is divisible by $q^{\lfloor a(H)/2\rfloor}.$ As $\ell_{1}$ and $\ell_{2}$ are primitive prime divisors of $\Phi_{m_{1}}$ and $\Phi_{m_{2}},$ where $m_{1}>m_{2}$ are the two largest integers such that $\Phi_{m_{i}}$ divides $|H|$, it follows from Lemma 2.5 that $\ell_{1}\geq m_{1}+1$ and $\ell_{2}\geq m_{2}+1,$ in particular, $\ell_{1},\ell_{2}\geq 11$. So $(\ell_{1}\ell_{2},\chi(1))=1,$ and then by part $(i)$ of Lemmas 3.13.5, we obtain that $\chi(1)=\frac{1}{t}q^{c(H)}\alpha,$ where $1\leq t\leq 6$ is an integer, $c(H)=4,7,7,11,16$, respectively and $\alpha$ is an integer coprime to $p$. However, this is a contradiction since $\chi(1)$ is divisible by $q^{\lfloor a(H)/2\rfloor}>q^{c(H)}$. Thus $k=1$ and $r=p.$
5.4. Eliminating finite simple groups of Lie type in characteristic $p$.
We will show that if $G^{\prime}/M\cong S$ is a simple group of Lie type in characteristic $p$ and $S\neq{}^{2}{\rm F}_{4}(2)^{\prime},$ then $S\cong H.$ We will prove this by eliminating other possibilities of $S.$ Assume that $S$ is a simple group of Lie type in characteristic $p$ and $S$ is not the Tits group. We have shown that $G^{\prime}/M\cong S$ and $|S|_{p}=|H|_{p}=q^{a(H)}.$ Observe that if $\theta\in{\mathrm{Irr}}(S)$ is extendible to ${\mathop{\rm Aut}\nolimits}(S),$ then $\theta$ extends to $G/C,$ where $C/M=C_{G/M}(G^{\prime}/M),$ so $\theta(1)\in{\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G).$ In fact, we will choose $\theta$ to be a unipotent character of $S,$ so by results of Lusztig, $\theta$ is extendible to ${\mathop{\rm Aut}\nolimits}(S)$ apart from some exceptions (see [17, Theorem $2.5$]). We refer to [4, $13.8,13.9$] for the classification of unipotent characters and the notion of symbols. In Table 2, for each simple group of Lie type $S$ in characteristic $p,$ we list the $p$-part of some unipotent character of $S$ that is extendible to ${\mathop{\rm Aut}\nolimits}(S).$
Subcase $H\cong\textrm{F}_{4}(q),q>2.$
$(a)$ Case $S\cong{\mathrm{L}}_{n}^{\epsilon}(p^{b}),$ with $n\geq 2.$ We have $bn(n-1)=48f.$ If $n=2$ then $b=24f$ so $S={\mathrm{L}}_{2}(q^{24})$ and hence $S$ has a character of degree $q^{24}+1.$ Obviously this degree divides no degree of $G$ since $q^{24}+1\nmid|{\rm F}_{4}(q)|.$ Next if $n=3$ then $b=8f$ and so $S={\mathrm{L}}_{3}^{\epsilon}(q^{8}).$ By [4, $(13.8)$], $S$ possesses a unipotent character parametrised by the partition $(1,2)$ of degree $q^{8}(q^{8}+\epsilon 1).$ As $q^{8}+1$ does not divide the order of $H$ so the case when $\epsilon=+$ cannot happen. Now assume $\epsilon=-.$ Then $q^{8}(q^{8}+\epsilon 1)=q^{8}(q^{8}-1)=q^{8}\Phi_{1}\Phi_{2}\Phi_{4}\Phi_{8},$ which is coprime to $\ell_{3}\ell_{4},$ so $q^{8}(q^{8}-1)$ must be one of the degrees given in case $(iv)$ of Lemma 3.1. However by comparing the power of $q,$ we see that this case is impossible. If $n=4,$ then $b=4f$ so $S={\mathrm{L}}_{4}^{\epsilon}(q^{4}).$ In this case, the unipotent character parametrized by the partition $(2,2)$ has degree $q^{8}(q^{8}+1).$ As above, this degree does not belong to ${\mathrm{cd}}(G).$ Thus we can assume that $n\geq 5.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)(n-2)/2}.$ By Lemma 3.1$(vi)$, we have $b(n-1)(n-2)/2\leq 16f$ if $p$ is odd and $b(n-1)(n-2)/2\leq 13f$ if $p$ is even. Assume first that $p$ is even. Multiplying both sides of the inequality $b(n-1)(n-2)/2\leq 13f$ by $n,$ we obtain $bn(n-1)(n-2)/2=24f(n-2)\leq 13nf,$ and so $24(n-2)\leq 13n.$ Thus $11n\leq 48$ so $n\leq 4,$ which is a contradiction. Thus $p$ is odd. Using the same argument with $b(n-1)(n-2)/2\leq 16f,$ we obtain that $n\leq 6.$ Hence $5\leq n\leq 6$ and $p$ is odd. Assume that $n=5.$ Then $5b=12f.$ We have that the primitive prime divisor $\ell_{4b}(p)$ exists and so as $\ell_{4b}(p)\mid|S|,$ we deduce that $\ell_{4b}(p)\mid|H|.$ Since $4b=48f/5>8f,$ we must have that $4b\mid 12f=5b,$ which is impossible. Now assume $n=6.$ Then $5b=8f$ and the primitive prime divisor $\ell_{6b}(p)\in\pi(S)\subseteq\pi(H)$ exists. As $6b>5b=8f$ we have that $6b\mid 12f,$ and hence $b\mid 2f,$ which implies that $5b=8f\mid 10f,$ which is impossible.
[…]
If $n=4$ then $2b=3f$ so $b\geq 3.$ Hence $\ell_{6b}(p)\in\pi(S)\subseteq\pi(H),$ and $6b=9f>8f,$ we deduce that $6b=9f\mid 12f,$ which is impossible.
If $n=5$ then $25b=24f.$ We have that $10b=240f/25=48f/5>9f$ so $\ell_{10b}(p)\in\pi(S)\subseteq\pi(H),$ hence $10b\mid 12f,$ which implies that $5b\mid 6f$ and so $25b=24f\mid 30f,$ which is impossible.
Hence we can assume that $n\geq 6.$ By Table 2, $S$ possesses a nontrivial irreducible character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)^{2}}/(2,p).$ Assume first that $p=2.$ Then $\chi(1)_{p}=p^{b(n-1)^{2}-1}$ and hence by 3.1$(vi),$ we obtain $bn(n-2)\leq b(n-1)^{2}-1\leq 13f.$ Multiplying both sides by $n$ and simplifying, we have $11n\leq 48$ so $n<5,$ which is a contradiction. Now assume $p>2.$ Then $\chi(1)_{p}=p^{b(n-1)^{2}}$ so by Lemma 3.1$(vi),$ we have that $b(n-1)^{2}\leq 16f.$ Multiplying both sides by $3,$ we obtain $3b(n-1)^{2}\leq 48f=2bn^{2}.$ After simplifying, we have $n(n-6)+3\leq 0,$ which is impossible as $n\geq 6.$
$(c)$ Case $S\cong{\mathrm{O}}_{2n}^{\epsilon}(q_{1}),$ where $q_{1}=p^{b},$ and $n\geq 4.$ We have $bn(n-1)=24.$ If $n=4,$ then $b=2f$ hence $q_{1}=q^{2}.$ Now if $S={\mathrm{O}}_{8}^{+}(q^{2}),$ then $S$ has a unipotent character $\chi$ labeled by the symbol in Table 2 with $\chi(1)=q^{14}\Phi_{8}^{2}$ and if $S={\mathrm{O}}_{8}^{-}(q^{2}),$ then $S$ has a unipotent character $\chi$ labeled by the symbol $\binom{1~{}3}{~{}-}$ with degree $\chi(1)={(q^{8}+1)(q^{6}-q^{2})}/{(q^{4}-1)}.$ However these degrees do not divide the order of $H.$ If $n=5,$ then $5b=6f.$ We have that $8b=48f/5>9f$ and so $\ell_{8b}(p)\in\pi(S)\subseteq\pi(H),$ so $8b\mid 12f=10b,$ which is impossible. If $n=6,$ then $5b=4f.$ We have that $8b=32f/5>6f$ and so $\ell_{8b}(p)\in\pi(S)\subseteq\pi(H),$ so $8b\mid 8f=10b$ or $8b\mid 12f=15b.$ However both cases are impossible. Thus we can assume that $n\geq 7.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{b(n-1)(n-2)}.$ By Lemma 3.1$(vi),$ we have $b(n-1)(n-2)\leq 16f.$ Multiplying both sides by $n,$ and simplifying, we obtain $8n\leq 48$ and hence $n\leq 6,$ which is a contradiction.
$(d)$ Case $S\cong{\rm F}_{4}(p^{b}),$ where $b\geq 1.$ Then $24b=24f,$ hence $b=f$ so $S\cong H.$
$(e)$ Case $S\cong{\rm G}_{2}(q_{1}),$ where $q_{1}=p^{b}$. We have $6b=24f$ and so $b=4f.$ Thus $S={\rm G}_{2}(q^{4}),$ so $S$ has an irreducible character of degree $q^{24}-1$ by [5]. However this degree divides no degrees of $H.$
$(f)$ Case $S\cong{}^{3}{\rm D}_{4}(p^{b})$. Then $12b=24f$ and hence $b=2f.$ By [4, 13.9], $S$ possesses a unipotent character $\chi$ labeled by the symbol $\phi_{1,3}^{\prime\prime}$ with degree $\chi(1)=q^{14}\Phi_{24}.$ But then this degree does not divide the order of $H.$
$(g)$ Case $S\in\{{}^{2}{\rm B}_{2}(q_{1}^{2}),{}^{2}{\rm G}_{2}(q_{1}^{2}),{}^{2}{\rm F}_{4}(q_{1}^{2})\},$ where $q_{1}^{2}=2^{2n+1},3^{2n+1},2^{2n+1}$ with $n\geq 1,$ respectively. For each case, the equations $|S|_{p}=p^{24f}$ is impossible.
$(g)$ For the remaining cases, we can argue as follows. We have that $|S|_{p}=p^{24f}.$ Lemma 3.1$(vi)$ yields that $\chi(1)_{p}\leq p^{16f},$ where $\chi$ is the unipotent character listed in Table 2. Using these two properties, we will obtain a contradiction. For instance, assume that $S\cong{\rm E}_{8}(p^{b}),$ where $b\geq 1.$ Then $120b=24f$ and hence $5b=f.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{91b}.$ By Lemma 3.1$(vi),$ we have $91b\leq 16f=80b,$ a contradiction.
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$
$(a)$ Case $S\cong{\mathrm{L}}_{n}^{\epsilon}(p^{b}),$ with $n\geq 2.$ We have $bn(n-1)=72f.$ If $n=2$ then $b=36f$ so $S={\mathrm{L}}_{2}(q^{36})$ and hence $S$ has a character of degree $q^{36}+1.$ However this is impossible as this number does not divide the order of $H.$ Hence we can assume that $n\geq 3.$ If $(n,q)=(6,2)$ or $(3,2),$ then obviously the equation $bn(n-1)=72f$ cannot happen. Thus $\ell_{\epsilon bn}(p)\in\pi(S)\subseteq\pi(H)$ always exists by Lemma 2.5, and so $bn\leq 18f.$ Multiplying both sides by $n-1$ and simplifying, we obtain that $n\geq 5.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)(n-2)/2}.$ By Lemma 3.2(vi) we have $b(n-1)(n-2)/2\leq 25f.$ Multiplying both sides of the latter inequality by $n$ and simplifying, we obtain that $n\leq 6.$ If $n=5,$ then $5b=18f$ and hence $4b=72f/5>14f$ so $\ell_{4p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $4b\mid 18f=5b,$ which is impossible. Similarly if $n=6,$ then $5b=12f$ and hence $6b=72f/5>14f$ so $\ell_{6p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $6b\mid 18f.$ Hence $12b\mid 36f=15b,$ a contradiction.
$(b)$ Case $S\cong{\mathrm{S}}_{2n}(q_{1}),$ or ${\mathrm{O}}_{2n+1}(q_{1}),$ with $q_{1}=p^{b},$ $n\geq 2$ and $S\neq{\mathrm{S}}_{4}(2).$ We have $bn^{2}=36f.$ If $p=2$ and $2nb=6,$ then $n=3,b=1$ and so $bn^{2}=9=36f,$ a contradiction. Thus $\ell_{2bn}(p)\in\pi(S)\subseteq\pi(H)$ exists. Hence $2bn\leq 18f.$ Multiplying both sides by $n$ and simplifying, we obtain $n\geq 4.$ By Table 2, $S$ possesses a nontrivial irreducible character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{bn(n-2)},$ and hence by Lemma 3.2$(vi)$, $bn(n-2)\leq 25f.$ Multiplying both sides by $n$ and simplifying, we have $11n\leq 72$ so $n\leq 6.$ If $n=4,$ then $4b=9f$ and hence $6b=54f/4>12f$ so $\ell_{6p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $6b\mid 18f=8b,$ which is impossible. Similarly if $n=5,$ then $25b=36f$ and hence $10b=360f/25>14f$ so $\ell_{10p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $10b\mid 18f.$ It follows that $20b\mid 36f=25b,$ which is impossible. Thus $n=6$ and so $b=f.$ By [4], $S$ possesses a unipotent character $\chi^{\alpha}$ labeled by the symbol $\alpha=\binom{0~{}1~{}6}{~{}-~{}}$ with degree $\chi^{\alpha}(1)=(q^{6}-1)(q^{6}-q)/(2(q+1)).$ Obviously $\chi^{\alpha}(1)$ is nontrivial and different from $|H|_{p}$ and furthermore it is coprime to $\ell_{i}$ for $i=1,2,3,$ which contradicts Lemma 3.2$(i)$.
$(c)$ Case $S\cong{\mathrm{O}}_{2n}^{\epsilon}(q_{1}),$ where $q_{1}=p^{b},$ and $n\geq 4.$ We have $bn(n-1)=36f.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{b(n-1)(n-2)}.$ By Lemma 3.2$(vi)$, we have $b(n-1)(n-2)\leq 25f.$ Multiplying both sides by $n,$ and simplifying, we obtain $n\leq 6.$ If $n=4,$ then $b=3f$ and hence $q_{1}=q^{3}.$ By [4], $S$ possesses a unipotent character $\varphi$ labeled by the symbol $\binom{1~{}n}{0~{}1}$ when $\epsilon=+$ and $\binom{0~{}1~{}n}{~{}~{}1~{}~{}}$ when $\epsilon=-$ of degree $(q_{1}^{2n}-q_{1}^{2})(q_{1}^{2}-1)=(q_{1}^{8}-q_{1}^{2})/(q_{1}^{2}-1)=q^{6}\Phi_{9}\Phi_{18}.$ It follows that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{2}\ell_{3})=1.$ By Lemma 3.2$(ii)$, $\varphi(1)$ must be one of the degrees in $(ii).$ However both cases are impossible by comparing the power on $q.$ If $n=5,$ then $5b=9f$ and hence $8b=72f/5>14f$ so $\ell_{8p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $8b\mid 18f=10b,$ which is impossible. Thus $n=6$ and so $5b=6f.$ As in case $n=4,$ $S$ possesses a unipotent character $\varphi$ of degree $(q_{1}^{12}-q_{1}^{2})/(q_{1}^{2}-1)=q_{1}^{2}(q_{1}^{10}-1)/(q_{1}^{2}-1)=q_{1}^{2}(q^{12}-1)/(q_{1}^{2}-1).$ It follows that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{1}\ell_{3})=1,$ contradicting Lemma 3.2$(iii)$.
$(d)$ Case $S\cong{\rm G}_{2}(p^{b}),$ where $b\geq 1.$ We have $6b=36f.$ By [5], $S$ possesses an irreducible character of degree $p^{6b}-1=q^{36}-1.$ However $q^{36}-1$ does not divide $|H|,$ a contradiction.
$(e)$ Case $S\in\{{}^{2}{\rm B}_{2}(q_{1}^{2}),{}^{2}{\rm G}_{2}(q_{1}^{2}),$ where $q_{1}^{2}=2^{2n+1},3^{2n+1},$ respectively with $n\geq 1.$ In these cases, the equation $|S|_{p}=|H|_{p}$ cannot happen.
$(f)$ Case $S\cong{}^{2}{\rm F}_{4}(q_{1}^{2}),$ where $q_{1}^{2}=2^{2n+1},$ with $n\geq 1.$ Then $12(2n+1)=36f.$ We have that $\ell_{12(2n+1)}(2)\in\pi(S)\subseteq\pi(H),$ which implies that $12(2n+1)=36f\leq 18f,$ which is impossible.
$(h)$ Case $S\cong{\rm F}_{4}(p^{b})$. Then $24b=36f$ and so $2b=3f.$ By [4, 13.9], $S$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{9,2}$ with $\varphi(1)=q_{1}^{2}\Phi_{3}^{2}(q_{1})\Phi_{6}^{2}(q_{1})\Phi_{12}(q_{1})=q^{3}\Phi_{9}^{2}\Phi_{18}.$ Hence $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{1}\ell_{3})=1,$ contradicting Lemma 3.2$(iii)$.
$(i)$ Case $S\cong{}^{3}{\rm D}_{4}(p^{b})$. Then $12b=36f.$ We have $\ell_{12b}(p)\in\pi(S)$ but $\ell_{12b}(p)=\ell_{36f}(p)\not\in\pi(H).$
$(j)$ Case $S\cong{}^{2}{\rm E}_{6}(p^{b})$. Then $36b=36f$ and hence $b=f$ so $S\cong H.$
$(k)$ Case $S\cong{\rm E}_{7}(p^{b})$. Then $63b=36f$ and hence $7b=4f.$ By Table 1, $S$ possesses a unipotent character $\varphi\neq{\mathrm{St}}_{S}$ with $\varphi(1)_{p}=p^{46b}.$ By Lemma 3.2$(vi)$, we obtain that $46b\leq 25f,$ which implies that $184b=4\cdot 46b\leq 100f=175b,$ which is impossible.
$(l)$ Case $S\cong{\rm E}_{8}(p^{b})$. Then $120b=36f$ and hence $10b=3f.$ By Table 1 and Lemma 3.2$(vi)$, we have that $91b\leq 25f.$ It follows that $273b\leq 25\cdot 3f=250b,$ a contradiction.
$(m)$ Case $S\cong{\rm E}_{6}(p^{b})$. Then $36b=36f$ and hence $b=f.$ By [4, 13.9], $S$ has a unipotent character $\varphi$ labeled by the symbol $\phi_{6,1}$ with degree $q\Phi_{8}\Phi_{9}.$ We have that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,|H|_{p}\}$ is coprime to $\ell_{1}\ell_{2}\ell_{4},$ which contradicts Lemma 3.2$(i)$.
The arguments for the remaining simple groups ${\rm E}_{6}(q),{\rm E}_{7}(q)$ and ${\rm E}_{8}(q)$ are similar, so we will skip the details.
6. Verifying Step 3: $G^{\prime}$ is isomorphic to $H$
By Step $2,$ we obtain that $G^{\prime}/M\cong H.$ Hence it suffices to show that $M=1.$ By way of contradiction, assume that $M$ is nontrivial. Let $K\leq M$ be a normal subgroup of $G^{\prime}$ such that $M/K$ is a chief factor of $G^{\prime}.$ Assume first that $M/K$ is nonabelian. It follows that $M/K\cong S^{k},$ where $k\geq 1$ and $S$ is a nonabelian finite simple group. By Lemma 2.1 and [2, Lemma 5], $M/K$ possesses a nontrivial irreducible character $\varphi\in{\mathrm{Irr}}(M/K)$ which extends to $\varphi_{0}\in{\mathrm{Irr}}(G^{\prime}).$ As $G^{\prime}/M\cong H,$ $G^{\prime}/M$ possesses an irreducible character $\chi$ of degree $|H|_{p},$ and hence by Gallagher’s Theorem, $\varphi_{0}\chi$ is an irreducible character of $G^{\prime}$ of degree $\varphi_{0}(1)|H|_{p}>|H|_{p}.$ As $G^{\prime}\unlhd G,$ we obtain that $\varphi_{0}(1)|H|_{p}$ must divide some degree of $G,$ which contradicts [25, Lemma 2.4]. Thus $M/K$ must be a minimal normal elementary abelian subgroup of $G^{\prime}/K.$ Let $C/K$ be the centralizer of $M/K$ in $G^{\prime}/K.$ As $M/K$ is a minimal normal elementary abelian subgroup of $G^{\prime}/K,$ we deduce that $M/K\leq C/K\unlhd G^{\prime}/K.$ In particular, $M\unlhd C\unlhd G^{\prime}.$ As $G^{\prime}/M$ is nonabelian simple, either $C=G^{\prime}$ or $C=M.$
Assume $C=M.$ We have that $M/K$ is a self-centralizing minimal normal elementary abelian subgroup of $G^{\prime}/K$ and $G^{\prime}/M\cong H.$ Let $\chi\in{\mathrm{Irr}}(G^{\prime}/K)$ be such that $\chi(1)$ is the largest degree of $G^{\prime}/K.$ Applying Lemma 2.3 to $M/K\unlhd G^{\prime}/K,$ we have that $\chi(1)>b(H),$ which is a contradiction as $\chi(1)\in{\mathrm{cd}}(G^{\prime})$ must divide some degree of $H$ and thus $\chi(1)\leq b(H).$
Assume $C=G^{\prime}$. Then $M/K=Z(G^{\prime}/K).$ Since $G^{\prime}$ is perfect, we obtain that $G^{\prime}/K$ is perfect and so $M/K=(G^{\prime}/K)^{\prime}\cap Z(G^{\prime}/K),$ which implies that $G^{\prime}/K$ is a central extension of $G^{\prime}/M$ (see [10, p. 629]). We now consider each case separately.
Subcase $H\cong\textrm{F}_{4}(q),(q>2)$ or ${\rm E}_{8}(q)$. By [15, Theorem 5.1.4], the Schur multiplier of $G^{\prime}/M\cong H$ is trivial, so $G^{\prime}/K\cong G^{\prime}/M,$ which implies that $M/K$ is trivial, which is a contradiction.
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ By [15, Theorem 5.1.4], we know that the Schur multiplier of $G^{\prime}/M\cong H$ is cyclic of order $d=(3,q+1).$ If $d=1,$ then $G^{\prime}/K\cong H$ so $M/K$ is trivial, a contradiction. Hence $d=3.$ Assume first that $q>2.$ Then ${}^{2}{\rm E}_{6}(q)_{sc}$ is the universal covering group of $H$ so $G^{\prime}/K\cong{}^{2}{\rm E}_{6}(q)_{sc}.$ As $G^{\prime}\unlhd G$ and ${\mathrm{cd}}(G^{\prime}/K)\subseteq{\mathrm{cd}}(G^{\prime}),$ we deduce that every degree of $G^{\prime}/K$ must divide some degree of $G,$ which is impossible by Lemma 3.2$(x)$. Now assume that $q=2.$ We have that the Schur multiplier of ${}^{2}{\rm E}_{6}(2)$ is isomorphic to $\mathbb{Z}_{2}\times Z_{6}.$
If $G^{\prime}/K$ is isomorphic to $2\cdot{}^{2}{\rm E}_{6}(2)$ or $6\cdot{}^{2}{\rm E}_{6}(2),$ then $G^{\prime}/K$ possesses an irreducible character of degree $2^{7}\cdot 3^{5}\cdot 5^{2}\cdot 7^{2}\cdot 11\cdot 13\cdot 19$ which divides no degree of ${}^{2}{\rm E}_{6}(2)$ and if $G^{\prime}/K\cong 3\cdot{}^{2}{\rm E}_{6}(2),$ then $G^{\prime}/K$ has an irreducible character of degree $2^{11}\cdot 3^{3}\cdot 5\cdot 7^{2}\cdot 13\cdot 17\cdot 19$ Thus this case cannot happen.
The remaining cases can be argued similarly, we skip the details.
7. Verifying Step 4: $G=G^{\prime}\times C_{G}(G^{\prime}).$
We have shown that $G^{\prime}\cong H$. Let $C=C_{G}(G^{\prime}).$ Then $G/C$ is almost simple with socle $H.$ Assume that $G^{\prime}\times C<G.$ Then $G$ induces some outer automorphism on $G^{\prime}.$
We first recall the set up in Section 3. Let $\mathcal{G}$ be a simple simply connected algebraic group in characteristic $p$ and $F$ is a Frobenius map such that $\mathcal{G}^{F}\cong H_{sc}=L,$ where $H\cong L/Z(L)$ is one of the simple groups in (1). Denote by $\mathcal{G}^{*}$ the dual group of $\mathcal{G}$ and let $F^{*}$ be the corresponding Frobenius automorphism. We have that $H^{*}:=(\mathcal{G}^{*})^{F^{*}}\cong H_{ad}.$
Subcase $H\cong\textrm{F}_{4}(q),q>2.$ In this case, $H_{sc}\cong H\cong H_{ad}.$ Note that when $q$ is odd, $\textrm{Out}(H)$ is a cyclic group of order $f,$ consisting of field automorphisms, and when $p$ is even, $\textrm{Out}(H)$ is a cyclic group of order $2f$ with generator $\gamma$ such that $\gamma^{2}$ is a generator for the group of field automorphisms of $H.$
Assume that $\sigma$ is a nontrivial field automorphism of $H.$ Let $\ell=\ell_{12f}(p)$ be a primitive prime divisor of $p^{12f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(\mathcal{G}^{*})^{F^{*}}\cong H.$ We have that $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{12}(q)$ and $H$ possesses an irreducible character $\chi_{s}$ with degree $|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}.$ Let $\sigma^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\sigma.$ We obtain that $(\mathcal{G}^{*})^{\sigma^{*}}\cong{\rm F}_{4}(p^{b}),$ where $1\leq b<f,$ so that $\ell=\ell_{12f}(p)$ does not divide $|(\mathcal{G}^{*})^{\sigma^{*}}|,$ and hence by [6, Lemma 2.5(ii)], $\chi_{s}$ is not $\sigma$-invariant, and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ contradicting Lemma 3.1$(v)$.
Now assume that $\sigma$ is not a field automorphism of $H.$ Then $p=2$ and $\sigma=\gamma^{k}$ for some odd integer $1\leq k\leq f,$ where $\gamma$ is a generator of $\mathop{\rm Out}\nolimits(H).$ Let $\tau=\sigma^{2}=\gamma^{2k}.$ If $\tau$ is nontrivial, then it is a field automorphism of $H,$ by applying the same argument as above, we obtain that $\chi_{s}$ is not $\tau$-invariant and so obviously $\chi_{s}$ is not $\sigma$-invariant, which leads to a contradiction. Hence $\sigma^{2}=1,$ which forces $f=k$ is odd. Thus $\sigma$ is a graph automorphism of order $2$ of ${\rm F}_{4}(2^{f}),$ where $f\geq 2$ is odd. By [17, Theorem 2.5(e)], $H$ has two unipotent characters labeled by the symbol $\phi_{2,16}^{\prime}$ and $\phi_{2,16}^{\prime\prime}$ with degree $\frac{1}{2}q^{13}\Phi_{4}\Phi_{8}\Phi_{12},$ which are fused under $\sigma.$ Therefore $G/C$ possesses an irreducible character of degree $q^{13}\Phi_{4}\Phi_{8}\Phi_{12}.$ However this contradicts Lemma 3.1$(vi)$ as $q$ is even. Therefore, $G=G^{\prime}C=G^{\prime}\times C,$ since $C\cap G^{\prime}=1,$ and hence $C\cong G/G^{\prime}$ is abelian.
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ Note that $\textrm{Out}(H)$ is a group of order $2df,$ where $d=(3,q+1).$ Assume first that $G/C$ possesses only inner and diagonal automorphism. It follows that $G/C\cong{}^{2}{\rm E}_{6}(q)_{ad},$ $3\mid(q+1)$ and $\sigma$ is a diagonal automorphism of $H.$ Thus we obtain that ${\mathrm{cd}}({}^{2}{\rm E}_{6}(q)_{ad})={\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G)={\mathrm{cd}}(H),$ which is impossible by Lemma 3.2$(x)$.
Hence we can assume that $\sigma=\gamma\varrho,$ where $\varrho$ is a field automorphism and $\gamma$ is either trivial or a diagonal automorphism. Let $\ell=\ell_{18f}(p)$ be a primitive prime divisor of $p^{18f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(\mathcal{G}^{*})^{F^{*}}.$ We have that $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{18}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree $|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{}^{2}{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{18}$ which acts trivially on its center, so $\chi_{s}\in{\mathrm{Irr}}(H).$ Let $\gamma^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\gamma.$ We obtain that $(\mathcal{G}^{*})^{\gamma^{*}}\cong{}^{2}{\rm E}_{6}(p^{b})_{ad},$ where $1\leq b<f,$ so $\ell=\ell_{18f}(p)$ does not divide $|(\mathcal{G}^{*})^{\gamma^{*}}|,$ and hence by [6, Lemma 2.5(ii)], we deduce that $\chi_{s}$ is not $\gamma$-invariant. As $\chi_{s}$ is invariant under the diagonal automorphisms, we deduce that $\chi_{s}$ is not $\sigma$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible. Thus, $G=G^{\prime}C=G^{\prime}\times C,$ since $C\cap G^{\prime}=1,$ and hence $C\cong G/G^{\prime}$ is abelian.
Subcase $H\cong\textrm{E}_{6}(q).$ Note that $\textrm{Out}(H)\cong\langle\gamma\rangle(\langle\varrho\rangle\times\langle\tau\rangle))$ is a group of order $2df,$ where $d=(3,q-1),$ $\gamma$ is a diagonal automorphism of order $d,$ $\varrho$ is a field automorphism of order $f$ and $\tau$ is a graph automorphism of order $2$ (see for example [7, Theorem 2.5.12]). Moreover $\tau$ inverts the cyclic group $\langle\gamma\rangle.$
Assume first that $G/C$ possesses only inner and diagonal automorphism. It follows that $G/C\cong{\rm E}_{6}(q)_{ad},$ $13\mid q-$ 1 and $\sigma$ is a diagonal automorphism of $H.$ Thus ${\mathrm{cd}}({\rm E}_{6}(q)_{ad})={\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G)={\mathrm{cd}}(H),$ which is impossible by Lemma 3.3$(x)$.
Hence we assume next that $\sigma=\gamma\varrho,$ where $\varrho$ is a nontrivial field automorphism and $\gamma$ is either trivial or a diagonal automorphism. Let $\ell=\ell_{12f}(p)$ be a primitive prime divisor of $p^{12f}-1$ and let $s$ be a semisimple element of order $\ell$ in $K\cong{\rm F}_{4}(q)$ where $K\leq(H^{*})^{\prime}$ is chosen so that it is fixed under the involutory graph $\tau$ (see [7, Proposition 4.9.2]). We have that $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{12}\Phi_{3}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree $|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{12}\Phi_{3},$ which is trivial on the center as $s\in(H^{*})^{\prime}$ so $\chi_{s}\in{\mathrm{Irr}}(H).$ Let $\gamma^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\gamma.$ We obtain that $(\mathcal{G}^{*})^{\gamma^{*}}\cong{\rm E}_{6}(p^{b})_{ad},$ where $1\leq b<f,$ so $\ell=\ell_{12f}(p)$ does not divide $|(\mathcal{G}^{*})^{\gamma^{*}}|,$ and hence by [6, Lemma 2.5(ii)], we deduce that $\chi_{s}$ is not $\gamma$-invariant. As $\chi_{s}$ is invariant under the diagonal automorphisms, we deduce that $\chi_{s}$ is not $\sigma$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible.
Now assume that $\sigma=\gamma\varrho\tau,$ where $\gamma$ is either trivial or a diagonal automorphism, $\varphi$ is a nontrivial field automorphism and $\tau$ is a graph automorphism. In this case, as the field and graph automorphisms commute by [7, Theorem 2.5.12], we can write $\sigma=\gamma\tau\varrho.$ Choose the semisimple element $s\in(H^{*})^{\prime}$ as in previous paragraph with $|s|=\ell_{12}(q).$ Since $s$ in $\tau$-invariant, [22, Corollary 2.5] yields that $\chi_{s}^{\tau}=\chi_{\tau(s)}=\chi_{s}.$ As $\chi_{s}$ is also $\gamma$-invariant by applying [22, Corollay 2.5] again, $\chi_{s}$ is $\gamma\tau$-invariant. Thus $\chi_{s}^{\sigma}=\chi_{s}^{\gamma\tau\varrho}=\chi_{s}^{\varrho}\neq\chi_{s}.$ Therefore $\chi_{s}$ is not $\sigma$-invariant, which leads to a contradiction.
Finally assume that $\sigma=\gamma\tau,$ where $\gamma$ is either trivial or a diagonal automorphism and $\tau$ is a graph automorphism. Let $\ell=\ell_{9f}(p)$ be a primitive prime divisor of $p^{9f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(H^{*})^{\prime}.$ Let $T$ be a maximal torus of $H^{*}$ containing $s.$ It is well known that $T$ is uniquely determined up to conjugation in $H^{*}.$ We have that $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{9}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree $|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{9},$ which is trivial on the center as $s\in(H^{*})^{\prime}$ so $\chi_{s}\in{\mathrm{Irr}}(H).$ As $\chi_{s}$ is invariant under the diagonal automorphisms, it suffices to show that $\chi_{s}$ is not $\tau$-invariant so it is not $\sigma$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible. Now we know that ${\mathbf{C}}_{H^{*}}(\tau)\cong\textrm{F}_{4}(q)$ and thus $\ell\nmid|{\mathbf{C}}_{H^{*}}(\tau)|.$ Consider the $H^{*}$-conjugacy class $(s).$ Since $|s|=\ell,$ we deduce that the $H^{*}$-conjugacy class $(s)$ is not $\tau$-invariant. By [22, Corollary 2.5], $\chi_{s}$ is not $\tau$-invariant. Thus $G=G^{\prime}C=G^{\prime}\times C,$ since $C\cap G^{\prime}=1,$ and hence $C\cong G/G^{\prime}$ is abelian.
The arguments for the remaining cases are similar. The proof of the main theorem is now complete.