### v3

# Huppert’s conjecture for finite simple exceptional groups of Lie type

Abstract.
Let $G$ be a finite group and let ${\mathrm{cd}}(G)$ be the set of its complex irreducible character degrees. We show that if ${\mathrm{cd}}(G)={\mathrm{cd}}(H),$ where $H$ is a finite simple exceptional group of Lie type, then $G\cong H\times A,$ where $A$ is an abelian group. This completes the verification of Huppert’s Conjecture for all finite simple exceptional groups of Lie type.
Key words and phrases:
character degrees; exceptional groups of Lie type; Huppert’s Conjecture
2000 Mathematics Subject Classification:
Primary 20C15, secondary 20D05, 20G40
1. Introduction
Let $G$ be a finite group and let ${\mathrm{Irr}}(G)$ be the set of its complex irreducible characters. By a ‘character degree’ we mean the degree of a complex irreducible character. Let ${\mathrm{cd}}(G)$ be the set of all character degrees of $G$. One of the main questions in the character theory of finite groups is to determine which properties of a finite group can be determined by its character table. Note that the character degree set of a group can be read off from the character table. Thus it is interesting to ask how much the character degree set of a group knows about the group itself.
We cannot expect too much since the character degree set of a finite group $G$ and the character degree set of a direct product $G\times A$, where $A$ is an abelian group, coincide. Navarro [25] showed that the character degree set alone cannot determine the solvability of the group. He constructed a finite perfect group $H$ and a finite solvable group $G$ such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. Furthermore, Navarro and Rizo [27] found a finite perfect group and a finite nilpotent group with the same character degree set.
On the other hand, Huppert proposed in [12] that all finite nonabelian simple groups are essentially determined by the set of their character degrees: if $H$ is a finite nonabelian simple group and $G$ is a finite group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$, then $G\cong H\times A$ for some finite abelian group $A$. The main purpose of this paper is to prove the following theorem.
Theorem 1.1.
Let $G$ be a finite group and let ${H}$ be a finite simple exceptional group of Lie type. If ${\mathrm{cd}}(G)={\mathrm{cd}}({H}),$ then $G\cong H\times A,$ where $A$ is abelian.
Huppert’s Conjecture has been verified for various finite simple exceptional groups of Lie type (see [12, 29, 30, 32, 34]). To complete the verification of this conjecture for finite simple exceptional groups of Lie type, we verify the conjecture for the following finite simple groups: $\textrm{F}_{4}(q),~{}{}^{2}\textrm{E}_{6}(q),~{}\textrm{E}_{6}(q),~{}\textrm{E}_{7}(q),~{}\textrm{E}_{8}(q)$ where $q$ is a prime power.
Huppert’s Conjecture has been verified for all sporadic simple groups (see [31] and the references therein), the alternating groups [1] and many finite simple classical groups of small Lie rank (see [14]). The conjecture remains open for finite simple classical groups of large rank.
In [15], we extend the conjecture to include certain quasi-simple groups. Recently, an analog of Huppert’s Conjecture was proposed by Qian in [17, Problem 20.79] using codegrees. Recall that the codegree of a character $\chi\in{\mathrm{Irr}}(G)$ is the ratio $|G:{\mathrm{ker}}(\chi)|/\chi(1)$, where ${\mathrm{ker}}(\chi)$ denotes the kernel of $\chi$. Let $\textrm{cod}(G)$ be the set of all codegrees of $G$. Qian conjectured that if $G$ is a finite group and $H$ is a finite nonabelian simple group such that $\textrm{cod}(G)=\textrm{cod}(H)$, then $G\cong H.$ This conjecture has been verified for alternating groups, all sporadic simple groups, all finite simple exceptional groups of Lie type and finite simple projective special linear groups (see [28] and the references therein).
We now describe our approach to the proof of Huppert’s Conjecture. In [12], Huppert proposed a $5$-step strategy to verify his conjecture and this has been used successfully to verify the conjecture for many finite simple groups. In [1], we modified his method to verify the conjecture for alternating groups. We use this method and the data from [20] to prove the main theorem.
Suppose that $G$ is a finite group and $H$ is a finite nonabelian simple group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. To verify Huppert’s Conjecture for $H$, we need to prove the following.
Step 1: Show that $G^{\prime}=G^{\prime\prime}$.
Step 2: If $G^{\prime}/M$ is a chief factor of $G$, then $G^{\prime}/M\cong H$.
Step 3: Show that $M=1$.
Step 4: Show that $G\cong G^{\prime}\times C_{G}(G^{\prime})$.
To verify Steps $1$ and $2$, we follow the argument developed by Huppert. For Step 3, we can reduce to the case when $L$ is a quotient of $G^{\prime}$ and $N$ is a minimal normal elementary abelian subgroup of $L$ such that $L/N\cong H$ and either ${\mathbf{C}}_{L}(N)=N$ or ${\mathbf{C}}_{L}(N)=L$ (see the proof of Proposition 6.1). For the first possibility, we use a result due to Guralnick and Tiep [11] on the non-coprime $k(GV)$-problem to obtain a contradiction. If the latter holds, then $L$ is a quasi-simple group. This and Step 4 can now be dealt with using the data from [20].
The paper is organized as follows. We collect some results needed to verify Steps $1,2$ and $3$ in Section 2. In Section 3, we present all the information about the character degrees of the finite simple exceptional groups of Lie type that we need for the proof of the main theorem. We verify Steps 1–4 in Sections 47, respectively.
Notation.
If $n$ is an integer, then we denote by $\pi(n)$ the set of all prime divisors of $n.$ If $G$ is a group, then we write $\pi(G)$ to denote the set of all prime divisors of the order of $G.$ If $p$ is a prime and $n>1$ is an integer, then we write $n_{p}$ for the largest power of $p$ that divides $n$. Let $\rho(G)=\cup_{\chi\in{\mathrm{Irr}}(G)}\pi(\chi(1))$ be the set of all primes which divide some irreducible character degrees of $G.$ The largest character degree of $G$ is denoted by $b(G).$ For a group $G,$ let $k(G)$ be the number of conjugacy classes of $G.$ If $N\unlhd G$ and $\theta\in{\mathrm{Irr}}(N),$ then the inertia group of $\theta$ in $G$ is denoted by $I_{G}(\theta).$ The set of all irreducible constituents of $\theta^{G}$ is denoted by ${\mathrm{Irr}}(G|\theta).$ Finally, we denote by $\Phi_{n}=\Phi_{n}(q)$ the cyclotomic polynomial in variable $q.$ Other notation is standard.
2. Preliminaries
We collect various results that are needed for the proofs of the main theorem. We start with the following lemmas which will be used to verify Steps $2$ and $3$
Lemma 2.1.
If $S$ is a nonabelian simple group, then there exists a nontrivial irreducible character $\theta$ of $S$ that extends to ${\mathop{\rm Aut}\nolimits}(S).$ Moreover the following hold:
$(i)$ If $S$ is an alternating group of degree at least $7,$ then $S$ has two characters of consecutive degrees $n(n-3)/2$ and $(n-1)(n-2)/2$ that extend to ${\mathop{\rm Aut}\nolimits}(S).$
$(ii)$ If $S$ is a sporadic simple group or the Tits group, then $S$ has two nontrivial irreducible characters which extend to ${\mathop{\rm Aut}\nolimits}(S)$ and have coprime degrees. These characters are listed in Table 1,
$(iii)$ If $S$ is a simple group of Lie type then the Steinberg character ${\mathrm{St}}_{S}$ of $S$ of degree $|S|_{p}$ extends to ${\mathop{\rm Aut}\nolimits}(S).$
Proof.
This is essentially [2, Theorems $2,3,4$]. However for sporadic simple groups and the Tits group, we choose different pairs of coprime degrees from those in [2]. ∎
Lemma 2.1 will be used together with the following extension result.
Lemma 2.2.
Let $N$ be a minimal normal subgroup of a finite group $G$ such that $N\cong S^{k},$ where $S$ is a nonabelian simple group. If $\theta\in{\mathrm{Irr}}(S)$ extends to ${\rm{Aut}}(S),$ then $\theta^{k}\in{\mathrm{Irr}}(N)$ extends to $G.$
Proof.
This is Lemma $5$ in [2]. ∎
We also need the following number theoretic result. It is used to show that a certain character degree of the simple exceptional groups of Lie type cannot be written as a proper nontrivial power.
Lemma 2.3.
If $x,y,m\geq 2$ are integers and $x$ is a prime power, then the equation $x^{2}+x+1=y^{m}$ has no solution.
Proof.
The equation in the lemma can be rewritten as $(x^{3}-1)/(x-1)=y^{m}$. This is a special case of the Nagell-Ljunggren equation $(x^{n}-1)/(x-1)=y^{m}.$ Nagell and Ljunggren show (see [3]) that if $n=3$, then the only solution of the equation with $x,y,m\geq 2$ is $(x,y)=(18,7)$. However, $x=18$ is not a prime power. ∎
We next list the irreducible characters of prime power degrees of the finite nonabelian simple exceptional groups of Lie type. This is part of [23, Theorem 1.1].
Lemma 2.4.
Let $G$ be one of the finite simple exceptional groups of Lie type listed in (1) and let $\chi$ be a faithful irreducible character of $G$. Suppose that $\chi(1)=r^{d}$, where $r$ is a prime. Then $G$ is of characteristic $r$ and $\chi$ is the Steinberg character of $G$, so $\chi(1)=|G|_{r}$.
The next result is crucial to our proof of Step $3$ Unfortunately, we are unable to find a simple proof of this fact. The proof given below is an easy consequence of a deep result due to Guralnick and Tiep [11] concerning the non-coprime $k(GV)$-problem. For a finite group $G,$ let $b(G)$ be the largest character degree of $G.$ It follows from character theory that $|G|=\sum_{\chi\in{\mathrm{Irr}}(G)}\chi(1)^{2}$ and $k(G)=|{\mathrm{Irr}}(G)|,$ where $k(G)$ is the number of conjugacy classes of $G.$ Thus $|G|\leq k(G)b(G)^{2}$ and $b(G)^{2}\leq|G|.$
Lemma 2.5.
Let $L$ be a finite group and let $M$ be a minimal normal elementary abelian subgroup of $L.$ Assume that $L/M$ is isomorphic to a simple exceptional group of Lie type of Lie rank at least $4$ and that $C_{L}(M)=M.$ Then $b(L)>b(L/M).$
Proof.
Let $H\cong L/M$ be a finite simple exceptional group of Lie type of Lie rank at least $4$ By way of contradiction, assume that $b(L)\leq b(L/M)=b(H).$ Since $M$ is an abelian normal subgroup of $L$, the conjugation action of $L$ on $M$ induces an action of $H$ on $M$. Using this action, we form the semidirect product $HM$. By [11, Lemma 2.4], $k(L)\leq k(HM).$ Now
$|L|=\sum_{\chi\in{\mathrm{Irr}}(L)}\chi(1)^{2}\leq k(L)b(L)^{2}\leq k(HM)b(H)^{2}.$
As $H$ is a finite simple exceptional group of Lie type of Lie rank at least $4,$ and $H$ acts irreducibly on $M$. By [11, Theorems 1.4, 1.5], $k(HM)\leq|M|/2.$ Moreover, $|L|=|H|\cdot|M|$ and so the above inequality becomes
$|L|=|H|\cdot|M|\leq|M|b(H)^{2}/2.$
Hence $|H|\leq b(H)^{2}/2<b(H)^{2},$ a contradiction. Thus $b(L)>b(L/M)$ as wanted. ∎
To verify Step 3 for finite simple classical groups, we need the following result.
Problem 2.6.
Let $G$ be a finite perfect group and let $M$ be a minimal normal elementary abelian subgroup of $G$ such that $G/M$ is a finite simple classical group. If $C_{G}(M)=M$, then $G$ has an irreducible character $\chi$ such that $\chi(1)$ divides no character degree of $G/M.$
It is possible that $b(G)>b(G/M)$ in the above situation. To finish Step $3$, we need to show that if $L$ is a finite quasi-simple group with $L/Z(L)$ a finite simple classical group, then some character degree of $L$ divides no character degree of $L/Z(L)$.
Finally, we need the following result due to Zsigmondy.
Lemma 2.7.
([18, Theorems  5.2.14, 5.2.15]). Let $q$ and $n$ be integers with $q\geq 2,$ and $n\geq 3.$ Suppose that $(q,n)\neq(2,6).$ Then $q^{n}-1$ has a prime divisor $\ell$ such that $\ell\equiv 1$ (mod $n$) and $\ell$ does not divide $q^{k}-1$ for any $k<n.$ Moreover if $\ell\mid q^{k}-1$ for some integer $k,$ then $n\mid k.$
Such a prime $\ell$ is a primitive prime divisor (ppd for short) and is denoted by $\ell_{n}(q).$ If $n$ is odd and $(q,n)\neq(2,3)$ then $q^{2n}-1$ has a primitive prime divisor denoted by $\ell_{-n}(q).$ In terms of the cyclotomic polynomials, as $q^{n}-1=\prod_{k\mid n}\Phi_{k},$ if $\ell$ is a primitive prime divisor of $q^{n}-1,$ then $\ell\mid\Phi_{n}$ but $\ell\nmid\Phi_{k}$ for any $k<n.$ In this situation, we also say that $\ell$ is a primitive prime divisor of $\Phi_{n}.$
Table 1. Character degrees of sporadic simple groups and the Tits group
Group Character Degree Group Character Degree
$\textrm{M}_{11}$ $\chi_{8}$ $2^{2}\cdot 11$ O’N $\chi_{2}$ $2^{6}\cdot 3^{2}\cdot 19$
$\chi_{9}$ $3^{2}\cdot 5$ $\chi_{19}$ $7^{3}\cdot 11\cdot 31$
$\textrm{M}_{12}$ $\chi_{7}$ $2\cdot 3^{3}$ $\textrm{Co}_{3}$ $\chi_{3}$ $11\cdot 23$
$\chi_{8}$ $5\cdot 11$ $\chi_{6}$ $2^{7}\cdot 7$
$\textrm{J}_{1}$ $\chi_{5}$ $2^{2}\cdot 19$ $\textrm{Co}_{2}$ $\chi_{3}$ $11\cdot 23$
$\chi_{6}$ $7\cdot 11$ $\chi_{22}$ $3^{6}\cdot 5^{3}$
$\textrm{M}_{22}$ $\chi_{2}$ $3\cdot 7$ $\textrm{Fi}_{22}$ $\chi_{56}$ $2^{17}\cdot 11$
$\chi_{5}$ $5\cdot 11$ $\chi_{57}$ $3^{9}\cdot 7\cdot 13$
$\textrm{J}_{2}$ $\chi_{6}$ $2^{2}\cdot 3^{2}$ HN $\chi_{10}$ $3^{4}\cdot 11\cdot 19$
$\chi_{13}$ $5^{2}\cdot 7$ $\chi_{45}$ $2^{10}\cdot 5^{5}$
$\textrm{M}_{23}$ $\chi_{3}$ $3^{2}\cdot 5$ Ly $\chi_{7}$ $2^{8}\cdot 7\cdot 67$
$\chi_{9}$ $11\cdot 23$ $\chi_{50}$ $3\cdot 5^{6}\cdot 31\cdot 37$
HS $\chi_{2}$ $2\cdot 11$ Th $\chi_{2}$ $2^{3}\cdot 31$
$\chi_{7}$ $5^{2}\cdot 7$ $\chi_{7}$ $5^{3}\cdot 13\cdot 19$
$\textrm{J}_{3}$ $\chi_{6}$ $2^{2}\cdot 3^{4}$ $\textrm{Fi}_{23}$ $\chi_{4}$ $13\cdot 17\cdot 23$
$\chi_{13}$ $5\cdot 17\cdot 19$ $\chi_{94}$ $2^{18}\cdot 5^{2}\cdot 7\cdot 11$
$\textrm{M}_{24}$ $\chi_{7}$ $2^{2}\cdot 3^{2}\cdot 7$ $\textrm{Co}_{1}$ $\chi_{3}$ $13\cdot 23$
$\chi_{8}$ $11\cdot 23$ $\chi_{17}$ $2\cdot 5^{4}\cdot 7^{2}\cdot 11$
McL $\chi_{2}$ $2\cdot 11$ $\textrm{J}_{4}$ $\chi_{2}$ $31\cdot 43$
$\chi_{14}$ $3^{6}\cdot 7$ $\chi_{11}$ $2^{3}\cdot 3^{2}\cdot 23\cdot 29\cdot 37$
He $\chi_{9}$ $3\cdot 5^{2}\cdot 17$ $\textrm{Fi}_{24}^{\prime}$ $\chi_{2}$ $13\cdot 23\cdot 29$
$\chi_{15}$ $2^{7}\cdot 7^{2}$ $\chi_{6}$ $5^{2}\cdot 7^{3}\cdot 11\cdot 17$
Ru $\chi_{5}$ $3^{3}\cdot 29$ B $\chi_{2}$ $3\cdot 31\cdot 47$
$\chi_{20}$ $2^{2}\cdot 5^{3}\cdot 7\cdot 13$ $\chi_{119}$ $2^{39}\cdot 11\cdot 19\cdot 23$
Suz $\chi_{2}$ $11\cdot 13$ M $\chi_{2}$ $47\cdot 59\cdot 71$
$\chi_{43}$ $2^{10}\cdot 3^{5}$ $\chi_{16}$ $5^{9}\cdot 7^{6}\cdot 11^{2}\cdot 17\cdot 19$
${}^{2}\textrm{F}_{4}(2)^{\prime}$ $\chi_{8}$ $5^{2}\cdot 13$
$\chi_{20}$ $2^{6}\cdot 3^{3}$
3. Character degrees of simple exceptional groups of Lie type
We now collect the results concerning the character degrees of exceptional groups of Lie type that we need for the proof of Theorem 1.1. We first consider the following setup (see, for example, [22]). Let $\mathcal{G}$ be a simple simply connected algebraic group defined over the algebraic closure of a finite field ${\mathbb{F}}_{p}$ of characteristic $p$ and let $F:\mathcal{G}\longrightarrow\mathcal{G}$ be a Steinberg endomorphism such that the fixed point group $L=\mathcal{G}^{F}$ is finite and $H\cong L/Z(L)$ is a finite simple group. Let $\mathcal{G}^{*}$ be a group dual to $\mathcal{G}$ with a dual Frobenius map $F^{*}$ and let $H^{*}=({\mathcal{G}^{*}})^{F^{*}}.$ We use the notation $H_{sc}$ for $L$ and $H_{ad}$ for $H^{*}.$
According to Deligne - Lusztig theory (see [4, 7, 9]), there is a natural partition
${\mathrm{Irr}}(L)=\bigcup_{\begin{subarray}{c}s\in H^{*}/\sim\\ s\>\text{semisimple}\end{subarray}}\mathcal{E}(L,s)$
of ${\mathrm{Irr}}(L),$ where the union runs over $H^{*}$-conjugacy classes of semisimple elements. The irreducible characters of $L$ in $\mathcal{E}(L,1)$ are unipotent characters of $L.$ The unipotent character degrees of simple exceptional groups of Lie type are given explicitly in [4, 13.9]. Moreover for each $s,$ there is a bijection map
$\psi_{s}:\mathcal{E}(L,s)\longrightarrow\mathcal{E}({\mathbf{C}}_{H^{*}}(s),1)$
such that for each $\chi\in\mathcal{E}(L,s),$ the degree of $\chi$ is given by
$\chi(1)=|H^{*}:{\mathbf{C}}_{H^{*}}(s)|\psi_{s}(\chi)(1)$
(see for example [24]). The semisimple character corresponding to the conjugacy class with representative $s$ is the character $\chi_{s}$ such that $\psi_{s}(\chi_{s})$ is the principal character of ${\mathbf{C}}_{H^{*}}(s).$ In this case $\chi_{s}(1)=|H^{*}:{\mathbf{C}}_{H^{*}}(s)|.$
By results of Lusztig (see [21]), we know that all the unipotent characters of $L$ are trivial on the center of $L$; therefore, they are exactly all the unipotent characters of $H\cong L/Z(L).$ Moreover, unipotent characters of $H$ extend to $\mathop{\rm Aut}\nolimits(H),$ apart from some exceptions (see [21, Theorem 2.5]). This fact will be used frequently in the proofs of Steps $2$ and $3$ Some results on semisimple characters of $L$ can be found in [8, Lemma 2.5]. In Table 2, we list the $p$-parts of the degrees of some unipotent characters of finite simple groups of Lie type in characteristic $p$. Note that we follow the ATLAS notation [6] for finite simple groups of Lie type.
Table 2. Some unipotent characters of finite simple groups of Lie type
$S=S(p^{b})$ Symbol $p$-part of degree
$\textrm{L}_{n}^{\epsilon}(p^{b}),n\geq 3$ $(1^{n-2},2)$ $p^{b(n-1)(n-2)/2}$
[…]
$\textrm{S}_{2n}(p^{b}),p>2$ $p^{b(n-1)^{2}}$
[…]
${}^{3}\textrm{D}_{4}(p^{b})$ $\phi_{1,3}^{\prime\prime}$ $p^{7b}$
$\textrm{F}_{4}(p^{b})$ $\phi_{9,10}$ $p^{10b}$
${}^{2}\textrm{F}_{4}(q^{2})$ ${}^{2}B_{2}[a],\epsilon$ $\frac{1}{\sqrt{2}}q^{13}$
$\textrm{E}_{6}(p^{b})$ $\phi_{6,25}$ $p^{25b}$
${}^{2}\textrm{E}_{6}(p^{b})$ $\phi_{2,16}^{\prime\prime}$ $p^{25b}$
$\textrm{E}_{7}(p^{b})$ $\phi_{7,46}$ $p^{46b}$
$\textrm{E}_{8}(p^{b})$ $\phi_{8,91}$ $p^{91b}$
The results in this section are based on the data given in [20]. From this, we can extract all the information about the character degrees of both $L$ and $H^{*},$ the simply connected and adjoint versions of the simple exceptional group of Lie type $H.$ Note that if $d:=|Z(L)|=|H^{*}:H|$ is trivial, then all these versions of $H$ are isomorphic and we obtain the character degrees of the simple group $H.$ Otherwise, we only know that ${\mathrm{cd}}(H)\subseteq{\mathrm{cd}}(L)$ and every character degree of $H$ divides some degree of $H^{*}.$
We are now ready to list the results on the character degree sets of finite simple exceptional groups of Lie type. From now on we assume that $q=p^{f}$ is a power of a prime $p,$ where $f\geq 1$ and $H$ is isomorphic to one of the following simple groups:
$\textrm{F}_{4}(q),~{}{}^{2}\textrm{E}_{6}(q),~{}\textrm{E}_{6}(q),~{}\textrm{E}_{7}(q),~{}\textrm{E}_{8}(q).$ (1)
Since $H\cong\textrm{F}_{4}(2)$ has been handled in [33], we assume that $q\geq 3$ when $H\cong\textrm{F}_{4}(q).$ As a demonstration, we give a detailed proof for the case $H\cong{}^{2}{\rm E}_{6}(q)$. The remaining cases are similar and we skip their proofs.
Lemma 3.1.
Let $H$ be the simple exceptional group ${}^{2}{\rm E}_{6}(q)$ and let $\ell_{i},i=1,2,3,4$ be primitive prime divisors of $\Phi_{18},\Phi_{12},\Phi_{8}$ and $\Phi_{10},$ respectively. Let $x\neq q^{36}$ be a nontrivial character degree of $L\cong{}^{2}{\rm E}_{6}(q)_{sc}.$ Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbols ${}^{2}E_{6}[\theta^{i}],i=1,2$ with degree $\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}.$
  2. $(ii)$
    If $(\ell_{2}\ell_{3},x)=1,$ then $x$ is the degree of the unipotent characters labeled by the symbols $\phi_{8,3}^{\prime}$ or $\phi_{8,9}^{\prime\prime}$ with degree $\frac{1}{2}q^{3}\Phi_{2}^{4}\Phi_{6}^{2}\Phi_{10}\Phi_{18}~{}\mbox{or~{}}~{}\frac{1}{2}q^{15}\Phi_{2}^{4}\Phi_{6}^{2}\Phi_{10}\Phi_{18}.$
  3. $(iii)$
    We have $(\ell_{1}\ell_{3},x)>1$ and $(p\ell_{1}\ell_{4},x)>1.$
  4. $(iv)$
    If $q=2$ then $(p\ell_{3}\ell_{4},x)>1,$ and if $q>2$ and $(p\ell_{3}\ell_{4},x)=1,$ then $x=\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}.$
  5. $(v)$
    $H$ possesses two semisimple irreducible characters $\chi_{i},i=1,2$ corresponding to the maximal tori with orders $\Phi_{18}$ and $\Phi_{12}\Phi_{6},$ with degrees
    $\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{3}\Phi_{8}\Phi_{10}\Phi_{12},\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}\Phi_{10}\Phi_{18},$
    respectively, such that no proper multiple of each degree is a degree of $H.$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{25}.$
  7. $(vii)$
    The unipotent character labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $q\Phi_{8}\Phi_{18}$ is the smallest nontrivial character degree of $H.$
  8. $(viii)$
    Assume $x,y\in{\mathrm{cd}}(L)-\{1,q^{36}\}.$ If $p\neq 3,$ then $(x,y)>1.$ If $p=3$ and $(x,y)=1,$ then $\{x,y\}=\{\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10},~{}\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}\}$.
  9. $(ix)$
    $L$ has no consecutive degrees.
  10. $(x)$
    Suppose that $13\mid q+$ 1. There exists $\chi\in{\mathrm{Irr}}(L)$ of degree $\frac{1}{3}q^{9}\Phi_{1}^{3}\Phi_{2}^{4}\Phi_{3}\Phi_{4}^{2}\Phi_{6}\Phi_{8}\Phi_{10}\Phi_{12}$ with multiplicity $6$ such that $\chi(1)$ divides no degree of $H.$ Similarly $H^{*}\cong{}^{2}{\rm E}_{6}(q)_{ad}$ has an irreducible character $\psi$ with $\psi(1)=q^{9}\Phi_{1}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\Phi_{12}\Phi_{18}$ with multiplicity $1$ such that $\psi(1)$ is not a degree of $H.$
Proof.
The lists of character degrees of ${}^{2}{\rm E}_{6}(q)_{sc}$ and ${}^{2}{\rm E}_{6}(q)_{ad}$ are available in [20]. Statements $(i)-(vi)$ can be checked easily. For $(vii)$, see [19]. For $(viii),$ suppose that $x,y\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(x,y)=1.$ Assume first that $(\ell_{1}\ell_{2},x)=1.$ By $(i),$ $x=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}$ and so $p\ell_{3}\ell_{4}\mid x.$ Therefore, $(p\ell_{3}\ell_{4},y)=1.$ It follows from $(iv)$ that $q>2$ and $y=\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}.$ It suffices to show that $p=3.$ Assume that $p\neq 3.$ Then $q^{2}-1$ is divisible by $3$ Now $\Phi_{1}^{4}\Phi_{2}^{4}=(q^{2}-1)^{4}$ is divisible by $3^{4}$ so $x=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}$ is divisible by $3$ Also
$\Phi_{3}\Phi_{6}=q^{4}+q^{2}+1=(q^{4}-1)+(q^{2}-1)+3=(q^{2}-1)(q^{2}+2)+3$
is divisible by $3$ and thus $3$ divides both $x$ and $y$ and hence $(x,y)>1,$ a contradiction. Thus $p=3$ as required. Hence we can assume that $(\ell_{1}\ell_{2},x)>1.$ By interchanging the roles of $x$ and $y,$ we can also assume that $(\ell_{1}\ell_{2},y)>1.$ Without loss of generality, we assume that $\ell_{2}\mid x.$ So, $(\ell_{2},y)=1$ hence $\ell_{1}\mid y$ and thus $(\ell_{1},x)=1.$ By $(iii),$ $\ell_{3}\mid x$ so $(\ell_{3},y)=1$ and hence $(\ell_{2}\ell_{3},y)=1.$ By $(ii)$, $p\ell_{1}\ell_{4}\mid y$ and thus $(p\ell_{1}\ell_{4},x)=1,$ which contradicts $(iii)$. This proves claim $(viii)$.
For $(ix),$ without loss of generality, assume that $x=y+1.$ By $(vii),$ we can assume that $x>y>1.$ But $q^{36}\pm 1$ do not divide $|L|,$ so $L$ has no such degree; therefore, we assume that $x,y\in{\mathrm{cd}}(L)-\{1,q^{36}\}.$ Moreover $(x,y)=1$ and so by $(viii),$ $p=3$ and
$\{x,y\}=\{\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10},\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}\}.$
Since $q\geq 3,$
$\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}>q^{25}>q^{17}>\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}$
and thus $x=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}~{}\mbox{and}~{}y=\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}.$ But then
$y+1<q^{17}+1<q^{18}<q^{25}<x$
which is a contradiction as $x=y+1.$
For $(x),$ assume that $13\mid q+$ 1 . Using [20], ${}^{2}{\rm E}_{6}(q)_{sc}$ possesses an irreducible character $\chi$ with degree and multiplicity as in $(x).$ Also $\chi(1)$ divides no other degree of $L.$ Since ${\mathrm{cd}}(H)\subseteq{\mathrm{cd}}(L),$ $\chi(1)$ divides no degree of $H-\{\chi(1)\}.$ Hence it suffices to show that $\chi(1)$ is not a degree of $H.$ Note that $H$ is a normal subgroup of index $3$ in ${}^{2}{\rm E}_{6}(q)_{ad}$ and so, for every irreducible character $\varphi$ of $H,$ either $\varphi(1)$ or $3\varphi(1)$ is a degree of ${}^{2}{\rm E}_{6}(q)_{ad}$ depending on whether $\varphi$ is extendible to ${}^{2}{\rm E}_{6}(q)_{ad}$ or not. Therefore, to show that $\chi(1)$ is not a degree of $H,$ we only need to check that neither $\chi(1)$ nor $3\chi(1)$ is a degree of ${}^{2}{\rm E}_{6}(q)_{ad}.$ Using [20], we check that $\chi(1)$ satisfies these properties, which completes the proof of our first claim. The second claim follows easily using [20]. The proof is now complete. ∎
Lemma 3.2.
Let $H\cong{\rm F}_{4}(q),$ where $q\geq 3,$ and let $x\in{\mathrm{cd}}(H)$ with $x\not\in\{1,q^{24}\}$. Let $\ell_{i},i=1,2,\cdots,4,$ be primitive prime divisors of $\Phi_{12},\Phi_{8},\Phi_{6}$ and $\Phi_{3},$ respectively. Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2}.$
  2. $(ii)$
    If $(\ell_{1},x)=1,$ then $p\mid x$ or $x=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}.$
  3. $(iii)$
    If $(\ell_{2},x)=1,$ then $p\mid x$ or $x$ is $\Phi_{3}\Phi_{6}\Phi_{12}~{}(q~{}odd)$ or $\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{12}.$
  4. $(iv)$
    If $(\ell_{3}\ell_{4},x)=1,$ then $x=\frac{1}{2}q\Phi_{4}\Phi_{8}\Phi_{12},$ $q^{3}\Phi_{4}^{2}\Phi_{8}\Phi_{12},$ $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8},$ $q^{9}\Phi_{4}^{2}\Phi_{8}\Phi_{12},$ or $\frac{1}{2}q^{13}\Phi_{4}\Phi_{8}\Phi_{12}.$
  5. $(v)$
    There exist $\chi_{i}\in{\mathrm{Irr}}(H),i=1,2,$ with degrees $\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}$ and $\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{12},$ respectively, such that no proper multiple of each degree is a degree of $H.$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{16}$ when $q$ is odd and at most $\frac{1}{2}q^{13}$ when $q$ is even.
  7. $(vii)$
    The smallest nontrivial character degree of $H$ is $\Phi_{3}\Phi_{6}\Phi_{12}$ when $q$ is odd and $\frac{1}{2}q\Phi_{1}^{2}\Phi_{3}^{2}\Phi_{8}$ when $q$ is even.
  8. $(viii)$
    Assume $x,y\in{\mathrm{cd}}(H)-\{1,q^{24}\}.$ If $p\neq 3,$ then $(x,y)>1.$ If $p=3$ and $(x,y)=1,$ then $\{x,y\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$
  9. $(ix)$
    $H$ has no consecutive degrees.
Lemma 3.3.
Let $H$ be the simple exceptional group ${\rm E}_{6}(q)$ and let $\ell_{i},i=1,2,3,4$ be primitive prime divisors of $\Phi_{12},\Phi_{9},\Phi_{8}$ and $\Phi_{5},$ respectively. Let $x\neq q^{36}$ be a nontrivial character degree of $L.$ Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbols $E_{6}[\theta^{i}],i=1,2,$ with degree $\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{5}\Phi_{8}.$
  2. $(ii)$
    We have $(x,\ell_{2}\ell_{3})>1.$
  3. $(iii)$
    If $(x,\ell_{1}\ell_{3})=1,$ then $x$ is the degree of unipotent characters labeled by the symbols $(D_{4},1)$ or $(D_{4},\epsilon)$ with degrees $\frac{1}{2}q^{3}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}\Phi_{9}~{}\mbox{or~{}}\frac{1}{2}q^{15}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}\Phi_{9}.$
  4. $(iv)$
    We have $(x,p\ell_{2}\ell_{4})>1,$ and if $(x,p\ell_{3}\ell_{4})=1$ then $x=\Phi_{3}^{2}\Phi_{6}\Phi_{9}\Phi_{12}.$
  5. $(v)$
    $H$ possesses two irreducible characters $\chi_{i},i=1,2$ with degrees $\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{5}\Phi_{8}\Phi_{9}$ and $\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{3}^{3}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{5}\Phi_{8}\Phi_{12},$ respectively such that no proper multiples of each degree is a degree of $H.$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{25}.$
  7. $(vii)$
    The unipotent character labeled by the symbol $\phi_{6,1}$ with degree $q\Phi_{8}\Phi_{9}$ is the smallest nontrivial character degree of $H.$
  8. $(viii)$
    Assume $x,y\in{\mathrm{cd}}(L)-\{1,q^{36}\}.$ If $p\neq 3,$ then $(x,y)>1.$ If $p=3$ and $(x,y)=1,$ then $\{x,y\}=\{\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{5}\Phi_{8},~{}\Phi_{3}^{2}\Phi_{6}\Phi_{9}\Phi_{12}\}.$
  9. $(ix)$
    $L$ has no consecutive degrees.
  10. $(x)$
    Assume $13\mid q-$ 1 . There exists $\chi\in{\mathrm{Irr}}(L)$ of degree $\frac{1}{3}q^{9}\Phi_{1}^{4}\Phi_{2}^{3}\Phi_{3}\Phi_{4}^{2}\Phi_{5}\Phi_{6}\Phi_{8}\Phi_{12}$ with multiplicity $6$ such that $\chi(1)$ divides no degree of $H.$ Moreover ${\rm E}_{6}(q)_{ad}$ has an irreducible character $\psi$ of degree $q^{9}\Phi_{2}\Phi_{4}^{2}\Phi_{5}\Phi_{6}^{2}\Phi_{8}\Phi_{9}\Phi_{12}$ with multiplicity $1$ such that $\psi(1)$ is not a degree of $H.$
Lemma 3.4.
Let $H$ be the simple exceptional group ${\rm E}_{7}(q)$ and let $\ell_{i},i=1,2,\cdots,5$ be primitive prime divisors of $\Phi_{18},\Phi_{14},$ $\Phi_{12},\Phi_{9},$ and $\Phi_{7},$ respectively. Let $x$ be a nontrivial character degree of $L$ with $x\neq q^{63}.$ Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbol $E_{7}[\xi]$ or $E_{7}[-\xi]$ with degree $\frac{1}{2}q^{11}\Phi_{1}^{7}\Phi_{3}^{3}\Phi_{4}^{2}\Phi_{5}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{12}.$
  2. $(ii)$
    If $(\ell_{2}\ell_{3},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbols $(D_{4},\epsilon_{1})$ or $(D_{4},\epsilon_{2})$ with degree
    $\frac{1}{2}q^{4}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}\Phi_{7}\Phi_{9}\Phi_{10}\Phi_{18}\text{ or }\frac{1}{2}q^{25}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}\Phi_{7}\Phi_{9}\Phi_{10}\Phi_{18}.$
  3. $(iii)$
    If $(\ell_{1}\ell_{3},x)=1,$ then $\ell_{4}\ell_{5}\mid x$ or $x$ is the degree of unipotent characters labeled by $(E_{6}[\theta^{i}],1)$ or $(E_{6}[\theta^{i}],\epsilon),$ where $i=1,2,$ with degrees
    $\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{5}\Phi_{7}\Phi_{8}\Phi_{10}\Phi_{14}\text{ or }\frac{1}{3}q^{16}\Phi_{1}^{6}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{5}\Phi_{7}\Phi_{8}\Phi_{10}\Phi_{14}.$
  4. $(iv)$
    We have $(p\ell_{2}\ell_{5},x)>1,$ and if $(\ell_{4}\ell_{5},x)=1,$ then $x$ is the degree of the unipotent characters labeled by $\phi_{512,11}$ or $\phi_{512,12}$ with degree $\frac{1}{2}q^{11}\Phi_{2}^{7}\Phi_{4}^{2}\Phi_{6}^{3}\Phi_{8}\Phi_{10}\Phi_{12}\Phi_{14}\Phi_{18}.$
  5. $(v)$
    $H$ possesses two semisimple irreducible characters $\chi_{i},i=1,2$ corresponding to the maximal tori with orders $\Phi_{18}\Phi_{2}$ and $\Phi_{14}\Phi_{2},$ respectively, and having degrees
    $\Phi_{1}^{7}\Phi_{2}^{6}\Phi_{3}^{3}\Phi_{4}^{2}\Phi_{5}\Phi_{6}^{3}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{10}\Phi_{12}\Phi_{14}\text{ and }\Phi_{1}^{7}\Phi_{2}^{6}\Phi_{3}^{3}\Phi_{4}^{2}\Phi_{5}\Phi_{6}^{3}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{10}\Phi_{12}\Phi_{18},$
    respectively such that no proper multiples of these degrees is in ${\mathrm{cd}}(L).$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{46}.$
  7. $(vii)$
    The unipotent character labeled by the symbol $\phi_{7,1}$ with degree $q\Phi_{7}\Phi_{12}\Phi_{14}$ is the smallest nontrivial character degree of $H.$
  8. $(viii)$
    If $x,y\in{\mathrm{cd}}(L)-\{1,q^{63}\},$ then $(x,y)>1.$
  9. $(ix)$
    $L$ has no consecutive degrees.
  10. $(x)$
    If $q$ is odd and $q\equiv\epsilon$ (mod $4$), where $q\in\{\pm 1\},$ then $L$ has an irreducible character $\chi$ of degree $q^{14}\Phi_{1}^{3}\Phi_{2}^{2}\Phi_{3}^{2}\Phi_{5}\Phi_{6}^{2}\Phi_{7}\Phi_{9}\Phi_{10}\Phi_{12}\Phi_{14}\Phi_{18}$ with multiplicity ${(q-\epsilon)}/{4}$ such that $\chi(1)$ divides no degree of $H.$ Moreover $E_{7}(q)_{ad}$ has a nontrivial degree $\psi(1)$ which is not a degree of $H.$ In particular, if $q\equiv 1$ (mod 4), then we can choose
    $\psi(1)=q^{28}\Phi_{2}^{3}\Phi_{3}\Phi_{6}^{2}\Phi_{9}\Phi_{10}\Phi_{12}\Phi_{14}\Phi_{18}$
    and if $q\equiv-1$ (mod 4), then $\psi(1)=q^{28}\Phi_{1}^{3}\Phi_{3}^{2}\Phi_{5}\Phi_{6}\Phi_{7}\Phi_{9}\Phi_{12}\Phi_{18}.$
Lemma 3.5.
Let $H$ be the simple exceptional group ${\rm E}_{8}(q)$ and let $\ell_{i},i=1,2,\cdots,5,$ be primitive prime divisors of $\Phi_{30},\Phi_{24},\Phi_{20},\Phi_{15}$ and $\Phi_{14},$ respectively. Let $x$ be a nontrivial character degree of $H$ with $x\neq q^{120}.$ Then the following hold.
  1. $(i)$
    If $(\ell_{1}\ell_{2},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbol $E_{8}[-\theta]$ or $E_{8}[-\theta^{2}]$ with degree $\frac{1}{6}q^{16}\Phi_{1}^{8}\Phi_{2}^{6}\Phi_{3}^{2}\Phi_{4}^{4}\Phi_{5}^{2}\Phi_{7}\Phi_{8}^{2}\Phi_{9}\Phi_{10}^{2}\Phi_{12}\Phi_{14}\Phi_{15}\Phi_{20}.$
  2. $(ii)$
    If $(\ell_{2}\ell_{3},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbol $E_{8}[i]$ or $E_{8}[-i]$ with degree $\frac{1}{4}q^{16}\Phi_{1}^{8}\Phi_{2}^{8}\Phi_{3}^{4}\Phi_{5}^{2}\Phi_{6}^{4}\Phi_{7}\Phi_{9}\Phi_{10}^{2}\Phi_{14}\Phi_{15}\Phi_{18}\Phi_{30}.$
  3. $(iii)$
    If $(\ell_{1}\ell_{3},x)=1,$ then $x$ is the degree of unipotent characters labeled by the symbol $E_{8}[\zeta^{k}],k=1,2,\cdots,4$ with degree $\frac{1}{5}q^{16}\Phi_{1}^{8}\Phi_{2}^{8}\Phi_{3}^{4}\Phi_{4}^{4}\Phi_{6}^{4}\Phi_{7}\Phi_{8}^{2}\Phi_{9}\Phi_{12}^{2}\Phi_{14}\Phi_{18}\Phi_{24}$ or $x$ is the degree of unipotent characters labeled by the symbol $(D_{4},\phi_{1,0})$ or $(D_{4},\phi_{1,24})$ with degrees $\frac{1}{2}q^{3}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}^{2}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{14}\Phi_{15}\Phi_{24},~{}\frac{1}{2}q^{63}\Phi_{1}^{4}\Phi_{3}^{2}\Phi_{5}^{2}\Phi_{7}\Phi_{8}\Phi_{9}\Phi_{14}\Phi_{15}\Phi_{24},$ respectively.
  4. $(iv)$
    We have $(p\ell_{4}\ell_{5},x)>1,$ and $(p\ell_{2}\ell_{5},x)>1.$
  5. $(v)$
    $H$ has irreducible characters $\chi_{i},i=1,2,3$ with degrees $|H|/\Phi_{m_{i}},$ where $m_{i}$ is $30,24$ and $20,$ respectively. Moreover no proper multiples of $\chi_{i}(1),i=1,2,3$ is a degree of $H.$
  6. $(vi)$
    The $p$-part of $x$ is at most $q^{91}.$
  7. $(vii)$
    The unipotent character with degree $q\Phi_{4}^{2}\Phi_{8}\Phi_{12}\Phi_{20}\Phi_{24}$ labeled by the symbol $\phi_{8,1}$ is the smallest nontrivial character degree of $H.$
  8. $(viii)$
    If $x,y\in{\mathrm{cd}}(H)-\{1,q^{120}\},$ then $(x,y)>1.$
  9. $(ix)$
    $H$ has no consecutive degrees.
4. Verifying Step 1: $G^{\prime}=G^{\prime\prime}$
We verify Step $1$ in this section. Let $\chi$ be an irreducible character of a finite group $G$. Following [29], $\chi$ is isolated in $G$ if $\chi(1)$ is divisible by no proper nontrivial character degree of $G,$ and no proper multiple of $\chi(1)$ is a character degree of $G.$ In this situation, $\chi(1)$ is an isolated degree of $G.$
Proposition 4.1.
Let $G$ be a finite group and let $H$ be one of the finite simple groups in (1). Assume that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. Then $G^{\prime}=G^{\prime\prime}.$
Proof.
By way of contradiction, suppose that $G^{\prime}\neq G^{\prime\prime}.$ Then there exists a normal subgroup $N$ of $G$ such that $G/N$ is solvable and minimal with respect to being nonabelian. By [29, Lemma 2.3], $G/N$ is an $r$-group for some prime $r$ or $G/N$ is a Frobenius group.
Case $1$ $G/N$ is an $r$-group. Since $G/N$ is a nonabelian $r$-group, it has an irreducible character $\tau$ of degree $r^{b}$ for some integer $b\geq 1$. By Lemma 2.4, the only nontrivial prime power degree of $H$ is the degree of the Steinberg character of $H,$ which is $|H|_{p}.$ It follows that $\tau(1)=r^{b}=|H|_{p}$ and hence $r=p.$ By Thompson’s Theorem ([16, Corollary 12.2]), $H$ has a nontrivial irreducible character $\psi$ with $p\nmid\psi(1)$. By [12, Lemma $2$], $\psi_{N}\in{\mathrm{Irr}}(N)$ and hence by Gallagher’s Theorem ([16, Corollary 6.17]), $\tau\psi$ is an irreducible character of $G$ of degree $\tau(1)\psi(1)=|H|_{p}\psi(1).$ But this contradicts [29, Lemma 2.4].
Case $2$ $G/N$ is a Frobenius group with kernel $F/N,$ $|F/N|=r^{a},$ $1<m=|G:F|\in{\mathrm{cd}}(G)$ and $r^{a}\equiv 1(\mbox{mod $m$}).$ By [29, Lemma 2.4], ${\mathrm{St}}_{H}(1)=|H|_{p}$ is an isolated degree of $H$ so by [29, Lemma 2.3(b)(2)], either $m=|H|_{p}$ or $r=p.$ We consider each family of simple groups separately.
Subcase $H\cong{\rm F}_{4}(q).$ Recall that $q>2$ in this case. Assume first that $m=|H|_{p}.$ As $r\nmid m,$ $r\neq p.$ By [4, $13.9$], $H$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $\varphi(1)=\frac{1}{2}q\Phi_{4}\Phi_{8}\Phi_{12}.$ As no proper multiple of $m$ is a character of $G$ and $\varphi(1)\nmid m,$ we deduce from [29, Lemma 2.3(b)(1)] that $r^{a}\mid\varphi(1)^{2}.$ As $r\neq p,$ $r^{a}\mid\varphi(1)_{p^{\prime}}^{2}=\frac{1}{4}\Phi_{4}^{2}\Phi_{8}^{2}\Phi_{12}^{2}.$ Now
$\frac{1}{2}\Phi_{4}\Phi_{8}\Phi_{12}=\frac{1}{2}(q^{4}+1)(q^{6}+1)<q^{10}.$
Hence $r^{a}\leq\varphi(1)_{p^{\prime}}^{2}<q^{20}.$ Since $m\mid r^{a}-1,$ $m=q^{24}\leq r^{a}-1<q^{20}-1<q^{20},$ a contradiction.
Thus $r=p.$ Hence $m\in{\mathrm{cd}}(G)$ is coprime to $p$ and so $m\neq|H|_{p}.$ Let $\chi_{i}\in{\mathrm{Irr}}(H),i=1,2,$ be irreducible characters of $H$ obtained from Lemma 3.2$(v)$. For each $i,$ as no proper multiple of $\chi_{i}(1)$ is a degree of $H,$ [29, Lemma 2.3(b)(2)] implies that $m$ divides $\chi_{i}(1)$ and hence $m$ divides $\gcd(\chi_{1}(1),\chi_{2}(1)).$ Thus $m$ is coprime to $\ell_{1}\ell_{2}.$ By Lemma 3.2$(i)$, $m=\frac{1}{4}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{6}^{2},$ which is impossible as $m$ is coprime to $p.$
Subcase $H\cong{}^{2}{\rm E}_{6}(q).$ Assume first that $m=|H|_{p}.$ As $r\nmid m,$ $r\neq p.$ By Lemma 3.1$(vii)$, $H$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{2,4}^{\prime}$ with degree $\varphi(1)=q\Phi_{8}\Phi_{18}.$ As no proper multiple of $m$ is a character of $G$ and $\varphi(1)\nmid m,$ by [29, Lemma 2.3(b)(1)], $r^{a}\mid\varphi(1)^{2}.$ As $r\neq p,$ $r^{a}\mid\varphi(1)_{p^{\prime}}^{2}=\Phi_{8}^{2}\Phi_{18}^{2}.$ Now
$\Phi_{8}\Phi_{18}=q^{10}-q^{7}+q^{6}+q^{4}-q^{3}+1<q^{15}.$
As $r^{a}\mid\varphi(1)_{p^{\prime}}^{2},$ $r^{a}\leq\varphi(1)_{p^{\prime}}^{2}<q^{30}.$ Since $m\mid r^{a}-1,$
$m=q^{36}\leq r^{a}-1<q^{30}-1<q^{30},$
which is impossible.
Thus $r=p.$ Then $m\in{\mathrm{cd}}(G)-\{1,|H|_{p}\}$ is coprime to $p.$ By [29, Lemma 2.3(b)(2)] and Lemma 3.1$(v),$ $m$ divides both $\chi_{i}(1),i=1,2.$ As $\chi_{i}(1)$ is coprime to $\ell_{i}$ for $i=1,2,$ $m$ is coprime to $\ell_{1}\ell_{2}$. By Lemma 3.1$(i)$, $m=\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}$, which is impossible as $p\nmid m.$
The arguments for the remaining cases are similar. ∎
5. Verifying Step 2: $G^{\prime}/M$ is isomorphic to $H$
Let $H$ be one of the simple groups listed in (1) and let $G$ be a finite group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. By Proposition 4.1, $G^{\prime}=G^{\prime\prime}$. Let $M\leq G^{\prime}$ be a normal subgroup of $G$ such that $G^{\prime}/M$ is a chief factor of $G$. The main purpose of this section is to show that $G^{\prime}/M\cong H.$ Recall that $H$ is a finite simple exceptional group of Lie type defined over a field of size $q=p^{f}$, where $p$ is a prime.
Proposition 5.1.
Let $G$ be a finite group and let $H$ be one of the finite simple groups in (1). Let $M\leq G^{\prime}$ be a normal subgroup of $G$ such that $G^{\prime}/M$ is a chief factor of $G$. Then $G^{\prime}/M\cong H.$
Proof.
By Proposition 4.1, $G^{\prime}$ is perfect. Hence $G^{\prime}/M\cong S^{k}$, where $S$ is a nonabelian simple group and $k\geq 1$ is an integer. Using the classification of finite simple groups, we first show that $S$ must be a finite simple groups of Lie type in characteristic $r$ for some prime $r$ and then show that $r=p,k=1$ and $S\cong H$.
(1) Assume $S=\mathrm{A}_{n},n\geq 7.$ Let $\theta_{i},i=1,2,$ be irreducible characters of $S$ obtained from Lemma 2.1$(i)$ with $\theta_{1}(1)=n(n-3)/2$ and $\theta_{2}(1)=(n-1)(n-2)/2.$ Now both $\theta_{i}$ extend to ${\mathop{\rm Aut}\nolimits}(\mathrm{A}_{n})\cong\mathrm{S}_{n}$ and $\theta_{2}(1)=\theta_{1}(1)+1.$ By Lemma 2.2, $\theta_{i}^{k}\in{\mathrm{Irr}}(G^{\prime}/M)$ extends to $G/M,$ for $i=1,2.$ Hence $\theta_{i}(1)^{k}\in{\mathrm{cd}}(G)$ and $\theta_{i}(1)^{k}$ are coprime for $i=1,2.$ Note that $\mathrm{S}_{n}$ also has an irreducible character labelled by the partition $(n-1,1)$ which is irreducible upon restriction to $\mathrm{A}_{n}$. Thus $\mathrm{A}_{n}$ has an irreducible character $\theta$ of degree $n-1$ which is extendible to $\mathop{\rm Aut}\nolimits(\mathrm{A}_{n})\cong\mathrm{S}_{n}$. Hence $\theta(1)^{k}=(n-1)^{k}\in{\mathrm{cd}}(G)$.
We need the following observation:
$\Phi_{3}\Phi_{6}\Phi_{12}=q^{8}+q^{4}+1<\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}=\frac{1}{3}q^{4}(q^{2}-1)^{2}(q^{4}-1)(q^{8}-1).$
If $q=3^{f},$ where $f\geq 1,$ then $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}$ is divisible by $q^{8}-1=3^{8f}-1,$ which is in turn divisible by $3^{8}-1=2^{5}\cdot 5\cdot 41,$ and so $\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}$ is always divisible by $2^{5}\cdot 3\cdot 5\cdot 41$
Subcase $H\cong\textrm{F}_{4}(q),q>2.$ Assume first that $p\neq 3.$ By Lemma 3.2$(viii),$ one of the degrees $\theta_{i}(1)^{k},i=1,2,$ must be $q^{24}.$ However, $(n-1,n-2)=1$ and $(n,n-3)=(n,3)$ so $\theta_{i}(1)^{k}$ can never be a power of $p\neq 3.$ Thus $p=3.$ If $\theta_{j}(1)^{k}=q^{24}$ for some $j,$ then $j=1$ and hence $\theta_{1}(1)^{k}=(n(n-3)/2)^{k}=3^{24f},$ where $q=3^{f}.$ As $n\geq 7,$ $n(n-3)/2$ is a prime power if and only if $n=9,$ so that $\theta_{1}(1)^{k}=3^{3k}=3^{24f}.$ Now $G$ has a degree $(n-1)^{k}=8^{k}$. By Lemma 2.4, the only nontrivial prime power degree of $G$ is $|H|_{p}=q^{24}=3^{24f},$ whence $3^{24f}=8^{k}$, a contradiction. Lemma 3.2$(viii)$ yields
$\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$
By Lemma 3.2$(ix)$, $G$ has no consecutive degrees, so $k\geq 2.$ But then $\Phi_{3}\Phi_{6}\Phi_{12}$ is a proper nontrivial power, which is impossible by Lemma 2.3.
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ Assume first that $p\neq 3.$ By Lemma 3.1$(viii),$ one of the degrees $\theta_{i}(1)^{k},i=1,2,$ must be $q^{36}.$ However $(n-1,n-2)=1$ and $(n,n-3)=(n,3)$ so $\theta_{i}(1)^{k}$ can never be a power of $p\neq 3.$ Thus $p=3.$ If $\theta_{j}(1)^{k}=q^{36}$ for some $j,$ then $j=1$ and hence $\theta_{1}(1)^{k}=(n(n-3)/2)^{k}=3^{36f},$ where $q=3^{f}.$ As $n\geq 7,$ $n(n-3)/2$ is a prime power if and only if $n=9,$ so that $\theta_{1}(1)^{k}=3^{3k}=3^{36f}.$ Again, $G$ has a degree $(n-1)^{k}=8^{k}$. However, by Lemma 2.4, the only nontrivial prime power degree of $G$ is $q^{36}=3^{36f},$ so we obtain a contradiction. Thus by Lemma 3.1$(viii)$,
$\{\theta_{1}(1)^{k},\theta_{2}(1)^{k}\}=\{\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18},\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\}.$
By Lemma 3.1$(ix)$, $G$ has no consecutive degrees so $k\geq 2.$ But then $\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}$ is a proper nontrivial power. As $q=3^{f},$ $\Phi_{3}$ is coprime to $\Phi_{6}^{2}\Phi_{12}\Phi_{18}$ and so $\Phi_{3}$ is a nontrivial proper power, which is impossible by Lemma 2.3.
The arguments for the remaining groups $\textrm{E}_{6}(q),\textrm{E}_{7}(q)$ and $\textrm{E}_{8}(q)$ are similar.
(2) Suppose that $S$ is a sporadic simple group or the Tits group. By Table 1, $S$ possesses two nontrivial irreducible characters $\theta_{i},i=1,2,$ such that $\theta_{1}(1)$ and $\theta_{2}(1)$ are coprime and each $\theta_{i}$ extends to $\mathop{\rm Aut}\nolimits(S).$ By Lemma 2.2, for each $i,$ $\theta_{i}^{k}$ extends to $G,$ so $\theta_{i}(1)^{k}\in{\mathrm{cd}}(G).$ Observe that for each $i,$ $\theta_{i}(1)$ is not a prime power so $\theta_{i}(1)^{k}\neq|H|_{p}.$
Assume $H\cong{\rm F}_{4}(q)$. As $(\theta_{1}(1)^{k},\theta_{2}(1)^{k})=1,$ by Lemma 3.2$(viii)$, $q=3^{f}$ and
$\{\theta_{1}(1)^{k},\theta_{2}(1)^{k}\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$
In this situation, the larger of $\theta_{i}(1)$ and $\theta_{2}(1)$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
Assume $H\cong{}^{2}{\rm E}_{6}(q)$. As $(\theta_{1}(1)^{k},\theta_{2}(1)^{k})=1,$ Lemma 3.1$(viii)$ yields that $q=3^{f}$ and
$\{\theta_{1}(1)^{k},\theta_{2}(1)^{k}\}=\{\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18},\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\}.$
In this situation, one of the degrees $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
Assume $H\cong{\rm E}_{6}(q)$. As $(\theta_{1}(1)^{k},\theta_{2}(1)^{k})=1,$ Lemma 3.3$(viii)$ implies that $q=3^{f}$ and
$\{\theta_{1}(1)^{k},\theta_{2}(1)^{k}\}=\{\Phi_{3}^{2}\Phi_{6}\Phi_{9}\Phi_{12},\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{5}\Phi_{8}\}.$
In this situation, one of the degrees $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
Assume $H\cong{\rm E}_{7}(q)$ or ${\rm E}_{8}(q)$. Observe that both $\theta_{i}(1)$ are not prime powers so both $\theta_{1}(1)^{k}$ and $\theta_{2}(1)^{k}$ are pairwise coprime and different from $|H|_{p},$ which contradicts Lemma 3.4$(viii)$ and Lemma 3.5$(viii)$, respectively.
(3) Assume that $S$ is a finite simple group of Lie type in characteristic $r,$ with $S\neq{}^{2}{\rm F}_{4}(2)^{\prime}.$ Let $\theta$ be the Steinberg character of $S.$ Then $\theta(1)=|S|_{r}$ and $\theta$ extends to ${\mathop{\rm Aut}\nolimits}(S).$ By Lemma 2.2, $\theta^{k}\in{\mathrm{Irr}}(G^{\prime}/M)$ extends to $G/M,$ hence $\theta(1)^{k}=|S|_{r}^{k}\in{\mathrm{cd}}(G).$ Since $|H|_{p}$ is the unique nontrivial prime power character degree of $G$ by Lemma 2.4, $\theta(1)^{k}=|H|_{p}.$ In particular, $r=p.$
We next claim that $k=1$. By way of contradiction, assume $k\geq 2.$ Write $\theta(1)=q_{1}^{s}$ and $|H|_{p}=q^{a(H)}$, so $a(H)=24,36,36,63,120$, respectively (see [18, Table 5.1.B]). Then $q_{1}^{sk}=q^{a(H)}.$ Let $\psi=\tau\times\theta\times\cdots\times\theta\in{\mathrm{Irr}}(S^{k}),$ where $\tau\in{\mathrm{Irr}}(S)$ with $1<\tau(1)\neq\theta(1).$ Then $\psi(1)=\tau(1)q_{1}^{s(k-1)}\nmid q^{a(H)}$ and is nontrivial, so it divides some character degree of $G,$ which is different from $q^{a(H)}.$ By part $(vi)$ of Lemmas 3.13.5, $q_{1}^{s(k-1)}=q^{a(H)(k-1)/k}\leq q^{b(H)}$, where $b(H)=25,16,25,46,91$, respectively, and hence $a(H)(k-1)\leq kb(H),$ which implies that $k\leq k_{0}$ for some positive integer $k_{0}$. We check that $k\leq k_{0}$, where $k_{0}=3$ unless $H={\rm E}_{8}(q)$ where $k_{0}=4$.
Let $C$ be a normal subgroup of $G$ such that $C/M=C_{G/M}(G^{\prime}/M).$ Then $G^{\prime}C/C\cong S^{k}$ is a unique minimal normal subgroup of $G/C$, so $G/C$ embeds into ${\mathop{\rm Aut}\nolimits}(S)\wr S_{k},$ where $S_{k}$ is the symmetric group of degree $k.$ Let $B={\mathop{\rm Aut}\nolimits}(S)^{k}\cap G/C.$ Then $|G/C:B|\mid k!.$ Let $\psi=1\times\theta\times\cdots\times\theta\in{\mathrm{Irr}}(G^{\prime}C/C).$ Then $\psi$ extends to $\psi_{0}\in{\mathrm{Irr}}(B)$. Let $J=I_{G}(\psi_{0})$ and let $\mu\in{\mathrm{Irr}}(J|\psi_{0}).$ By [16, Theorem 6.11], $\chi=\mu^{G}\in{\mathrm{Irr}}(G).$ Then
$\chi(1)=e\psi_{0}(1)=e\psi(1)=eq^{a(H)(k-1)/k},$
where $e\mid k!.$ Since $k\leq 4,$ we deduce that $e\mid 4!=24,$ so $1<\chi(1)<q^{a(H)}.$ In addition $\chi(1)$ is divisible by $q^{\lfloor a(H)/2\rfloor}.$ As $\ell_{1}$ and $\ell_{2}$ are primitive prime divisors of $\Phi_{m_{1}}$ and $\Phi_{m_{2}},$ where $m_{1}>m_{2}$ are the two largest integers such that $\Phi_{m_{i}}$ divides $|H|$, it follows from Lemma 2.7 that $\ell_{1}\geq m_{1}+1$ and $\ell_{2}\geq m_{2}+1$. In particular, $\ell_{1},\ell_{2}\geq 11$. So $(\ell_{1}\ell_{2},\chi(1))=1.$ By part $(i)$ of Lemmas 3.13.5, $\chi(1)=\frac{1}{t}q^{c(H)}\alpha,$ where $t=3,4,3,2,6$ and $c(H)=7,4,7,11,16$, respectively and $\alpha$ is an integer coprime to $p$. However, this is a contradiction since $\chi(1)$ is divisible by $q^{\lfloor a(H)/2\rfloor}>q^{c(H)}$. Thus $k=1$ and $r=p.$
(4) Eliminating finite simple groups of Lie type in characteristic $p$. We show that if $G^{\prime}/M\cong S$ is a simple group of Lie type in characteristic $p$ and $S\neq{}^{2}{\rm F}_{4}(2)^{\prime},$ then $S\cong H.$ We prove this by eliminating other possibilities of $S.$ Assume that $S$ is a simple group of Lie type in characteristic $p$ and $S$ is not the Tits group. We have shown that $G^{\prime}/M\cong S$ and $|S|_{p}=|H|_{p}=q^{a(H)}.$ Observe that if $\theta\in{\mathrm{Irr}}(S)$ is extendible to ${\mathop{\rm Aut}\nolimits}(S),$ then $\theta$ extends to $G/C,$ where $C/M=C_{G/M}(G^{\prime}/M),$ so $\theta(1)\in{\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G).$ In fact, we choose $\theta$ to be a unipotent character of $S.$ By [21, Theorem $2.5$], $\theta$ is extendible to ${\mathop{\rm Aut}\nolimits}(S)$ apart from some exceptions. We refer to [4, $13.8,13.9$] for the classification of unipotent characters and the notion of symbols. In Table 2, for each simple group of Lie type $S$ in characteristic $p,$ we list the $p$-part of some unipotent character of $S$ that is extendible to ${\mathop{\rm Aut}\nolimits}(S).$
Subcase $H\cong\textrm{F}_{4}(q),q>2.$
$(a)$ Case $S\cong{\mathrm{L}}_{n}^{\epsilon}(p^{b}),$ with $n\geq 2.$ Here $bn(n-1)=48f.$ If $n=2$ then $b=24f$, so $S={\mathrm{L}}_{2}(q^{24})$. Hence $S$ has a character of degree $q^{24}+1.$ Obviously this degree divides no degree of $G$ since $q^{24}+1\nmid|{\rm F}_{4}(q)|.$ If $n=3$ then $b=8f$ and so $S={\mathrm{L}}_{3}^{\epsilon}(q^{8}).$ By [4, $(13.8)$], $S$ possesses a unipotent character parametrized by the partition $(1,2)$ of degree $q^{8}(q^{8}+\epsilon 1).$ Since $q^{8}+1$ does not divide the order of $H$, the case when $\epsilon=+$ cannot occur. Now assume $\epsilon=-.$ Then $q^{8}(q^{8}+\epsilon 1)=q^{8}(q^{8}-1)=q^{8}\Phi_{1}\Phi_{2}\Phi_{4}\Phi_{8},$ which is coprime to $\ell_{3}\ell_{4},$ so $q^{8}(q^{8}-1)$ must be one of the degrees given in case $(iv)$ of Lemma 3.2. However by comparing the power of $q,$ we see that this case is impossible. If $n=4,$ then $b=4f$ so $S={\mathrm{L}}_{4}^{\epsilon}(q^{4}).$ In this case, the unipotent character parametrized by the partition $(2,2)$ has degree $q^{8}(q^{8}+1).$ As above, this degree does not belong to ${\mathrm{cd}}(G).$ Thus we assume that $n\geq 5.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)(n-2)/2}.$ By Lemma 3.2$(vi)$, if $p$ is odd, then $b(n-1)(n-2)/2\leq 16f$; else $b(n-1)(n-2)/2\leq 13f$. Assume first that $p$ is even. Multiplying both sides of the inequality $b(n-1)(n-2)/2\leq 13f$ by $n,$ we obtain
$bn(n-1)(n-2)/2=24f(n-2)\leq 13nf,$
and so $24(n-2)\leq 13n.$ Thus $11n\leq 48$ so $n\leq 4,$ a contradiction. Thus $p$ is odd. Using the same argument with $b(n-1)(n-2)/2\leq 16f,$ we obtain that $n\leq 6.$ Hence $5\leq n\leq 6$ and $p$ is odd. Assume that $n=5.$ Then $5b=12f.$ Since the primitive prime divisor $\ell_{4b}(p)$ exists, $\ell_{4b}(p)\mid|S|,$ and hence $\ell_{4b}(p)\mid|H|.$ Since $4b=48f/5>8f$, $4b\mid 12f=5b,$ which is impossible. Assume $n=6.$ Then $5b=8f$ and the primitive prime divisor $\ell_{6b}(p)\in\pi(S)\subseteq\pi(H)$ exists. As $6b>5b=8f$, $6b\mid 12f,$ and hence $b\mid 2f$. It follows that $5b=8f\mid 10f,$ which is impossible.
[…]
${q_{1}^{3}(q_{1}^{6}-1)}/({q_{1}^{2}-1})={q^{8}(q^{16}-1)}/({q^{16/3}-1}).$
However $\chi(1)$ is not a degree of $G,$ since $\ell_{16}(q)\mid\chi(1)$ but $\ell_{16}(q)\nmid|H|.$
If $n=4$ then $2b=3f$ so $b\geq 3.$ Hence $\ell_{6b}(p)\in\pi(S)\subseteq\pi(H),$ and $6b=9f>8f,$ and we deduce that $6b=9f\mid 12f,$ which is impossible.
If $n=5$ then $25b=24f.$ Now $10b=240f/25=48f/5>9f$ so $\ell_{10b}(p)\in\pi(S)\subseteq\pi(H),$ hence $10b\mid 12f,$ which implies that $5b\mid 6f$ and so $25b=24f\mid 30f,$ which is impossible.
Assume $n\geq 6.$ By Table 2, $S$ possesses a nontrivial irreducible character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)^{2}}/(2,p).$ Assume that $p=2.$ Then $\chi(1)_{p}=p^{b(n-1)^{2}-1}$. By Lemma 3.2$(vi),$
$bn(n-2)\leq b(n-1)^{2}-1\leq 13f.$
Multiplying both sides by $n$ and simplifying, $11n\leq 48$ so $n<5,$ a contradiction. Now assume $p>2.$ Then $\chi(1)_{p}=p^{b(n-1)^{2}}$. By Lemma 3.2$(vi),$ $b(n-1)^{2}\leq 16f.$ Multiplying both sides by $3,$
$3b(n-1)^{2}\leq 48f=2bn^{2}.$
After simplifying, $n(n-6)+3\leq 0,$ which is a contradiction as $n\geq 6.$
$(c)$ Case $S\cong{\mathrm{O}}_{2n}^{\epsilon}(q_{1}),$ where $q_{1}=p^{b},$ and $n\geq 4.$ Here $bn(n-1)=24.$ If $n=4,$ then $b=2f$ hence $q_{1}=q^{2}.$ If $S={\mathrm{O}}_{8}^{+}(q^{2}),$ then $S$ has a unipotent character $\chi$ labeled by the symbol in Table 2 with $\chi(1)=q^{14}\Phi_{8}^{2}$. If $S={\mathrm{O}}_{8}^{-}(q^{2}),$ then $S$ has a unipotent character $\chi$ labeled by the symbol $\binom{1~{}3}{~{}-}$ with degree $\chi(1)={(q^{8}+1)(q^{6}-q^{2})}/{(q^{4}-1)}.$ However, these degrees do not divide $|H|.$ If $n=5,$ then $5b=6f.$ Since $8b=48f/5>9f$, $\ell_{8b}(p)\in\pi(S)\subseteq\pi(H),$ so $8b\mid 12f=10b,$ which is impossible. If $n=6,$ then $5b=4f.$ Now $8b=32f/5>6f$ and so $\ell_{8b}(p)\in\pi(S)\subseteq\pi(H),$ thus $8b\mid 8f=10b$ or $8b\mid 12f=15b.$ Both cases are impossible. Thus we can assume that $n\geq 7.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{b(n-1)(n-2)}.$ By Lemma 3.2$(vi),$ $b(n-1)(n-2)\leq 16f.$ Multiplying both sides by $n,$ and simplifying, $8n\leq 48$ and hence $n\leq 6,$ a contradiction.
$(d)$ Case $S\cong{\rm F}_{4}(p^{b}),$ where $b\geq 1.$ Then $24b=24f,$ hence $b=f$ so $S\cong H.$
$(e)$ Case $S\cong{\rm G}_{2}(q_{1}),$ where $q_{1}=p^{b}$. Here $6b=24f$ and so $b=4f.$ Thus $S={\rm G}_{2}(q^{4}),$ so $S$ has an irreducible character of degree $q^{24}-1$ by [5]. This degree divides no degrees of $H.$
$(f)$ Case $S\cong{}^{3}{\rm D}_{4}(p^{b})$. Then $12b=24f$ and hence $b=2f.$ By [4, 13.9], $S$ possesses a unipotent character $\chi$ labeled by the symbol $\phi_{1,3}^{\prime\prime}$ with degree $\chi(1)=q^{14}\Phi_{24}.$ But then this degree does not divide the order of $H.$
$(g)$ Case $S\in\{{}^{2}{\rm B}_{2}(q_{1}^{2}),{}^{2}{\rm G}_{2}(q_{1}^{2}),{}^{2}{\rm F}_{4}(q_{1}^{2})\},$ where $q_{1}^{2}=2^{2n+1},3^{2n+1},2^{2n+1}$ with $n\geq 1,$ respectively. For each case, the equations $|S|_{p}=p^{24f}$ is impossible.
$(h)$ For the remaining cases, we argue as follows. In all cases $|S|_{p}=p^{24f}.$ Lemma 3.2$(vi)$ yields that $\chi(1)_{p}\leq p^{16f},$ where $\chi$ is the unipotent character listed in Table 2. Using these two properties, we obtain a contradiction. For instance, assume that $S\cong{\rm E}_{8}(p^{b}),$ where $b\geq 1.$ Then $120b=24f$ and hence $5b=f.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{91b}.$ By Lemma 3.2$(vi),$ $91b\leq 16f=80b,$ a contradiction.
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$
$(a)$ Case $S\cong{\mathrm{L}}_{n}^{\epsilon}(p^{b}),$ with $n\geq 2.$ Here $bn(n-1)=72f.$ If $n=2$ then $b=36f$ so $S={\mathrm{L}}_{2}(q^{36})$ and hence $S$ has a character of degree $q^{36}+1.$ However this is impossible as this number does not divide the order of $H.$ Hence we assume that $n\geq 3.$ If $(n,q)=(6,2)$ or $(3,2),$ then $bn(n-1)\neq 72f$. Thus $\ell_{\epsilon bn}(p)\in\pi(S)\subseteq\pi(H)$ always exists by Lemma 2.7, and so $bn\leq 18f.$ Multiplying both sides by $n-1$ and simplifying, $n\geq 5.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}=p^{b(n-1)(n-2)/2}.$ By Lemma 3.1(vi), $b(n-1)(n-2)/2\leq 25f.$ Multiplying both sides of the latter inequality by $n$ and simplifying, $n\leq 6.$ If $n=5,$ then $5b=18f$ and hence $4b=72f/5>14f$ so $\ell_{4p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $4b\mid 18f=5b,$ which is impossible. Similarly if $n=6,$ then $5b=12f$ and hence $6b=72f/5>14f$ so $\ell_{6p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $6b\mid 18f.$ Hence $12b\mid 36f=15b,$ a contradiction.
$(b)$ Case $S\cong{\mathrm{S}}_{2n}(q_{1}),$ or ${\mathrm{O}}_{2n+1}(q_{1}),$ with $q_{1}=p^{b},$ $n\geq 2$ and $S\neq{\mathrm{S}}_{4}(2).$ Now $bn^{2}=36f.$ If $p=2$ and $2nb=6,$ then $n=3,b=1$ and so $bn^{2}=9=36f,$ a contradiction. Thus $\ell_{2bn}(p)\in\pi(S)\subseteq\pi(H)$ exists. Hence $2bn\leq 18f.$ Multiplying both sides by $n$ and simplifying, $n\geq 4.$ By Table 2, $S$ possesses a nontrivial irreducible character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{bn(n-2)}$. By Lemma 3.1$(vi)$, $bn(n-2)\leq 25f.$ Multiplying both sides by $n$ and simplifying, $11n\leq 72$ so $n\leq 6.$ If $n=4,$ then $4b=9f$. Since $6b=54f/4>12f$, $\ell_{6p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $6b\mid 18f=8b,$ which is impossible. Similarly if $n=5,$ then $25b=36f$. Since $10b=360f/25>14f$, $\ell_{10p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $10b\mid 18f.$ It follows that $20b\mid 36f=25b,$ which is impossible. Thus $n=6$ and so $b=f.$ By [4], $S$ possesses a unipotent character $\chi^{\alpha}$ labeled by the symbol $\alpha=\binom{0~{}1~{}6}{~{}-~{}}$ with degree $\chi^{\alpha}(1)=(q^{6}-1)(q^{6}-q)/(2(q+1)).$ Obviously $\chi^{\alpha}(1)$ is nontrivial and different from $|H|_{p}$ and furthermore it is coprime to $\ell_{i}$ for $i=1,2,3,$ contradicting Lemma 3.1$(i)$.
$(c)$ Case $S\cong{\mathrm{O}}_{2n}^{\epsilon}(q_{1}),$ where $q_{1}=p^{b},$ and $n\geq 4.$ Here $bn(n-1)=36f.$ By Table 2, $S$ possesses a unipotent character $\chi\neq{\mathrm{St}}_{S}$ with $\chi(1)_{p}\geq p^{b(n-1)(n-2)}.$ By Lemma 3.1$(vi)$, $b(n-1)(n-2)\leq 25f.$ Multiplying both sides by $n,$ and simplifying, $n\leq 6.$ If $n=4,$ then $b=3f$ and hence $q_{1}=q^{3}.$ By [4], $S$ possesses a unipotent character $\varphi$ labeled by the symbol $\binom{1~{}n}{0~{}1}$ when $\epsilon=+$ and $\binom{0~{}1~{}n}{~{}~{}1~{}~{}}$ when $\epsilon=-$ of degree
$(q_{1}^{2n}-q_{1}^{2})(q_{1}^{2}-1)=(q_{1}^{8}-q_{1}^{2})/(q_{1}^{2}-1)=q^{6}\Phi_{9}\Phi_{18}.$
It follows that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{2}\ell_{3})=1.$ By Lemma 3.1$(ii)$, $\varphi(1)$ is one of the degrees in $(ii).$ Both cases are impossible by comparing the power on $q.$ If $n=5,$ then $5b=9f$ and hence $8b=72f/5>14f$ so $\ell_{8p}(p)\in\pi(S)\subseteq\pi(H)$ exists and thus $8b\mid 18f=10b,$ which is impossible. Thus $n=6$ and so $5b=6f.$ As when $n=4,$ $S$ possesses a unipotent character $\varphi$ of degree
$(q_{1}^{12}-q_{1}^{2})/(q_{1}^{2}-1)=q_{1}^{2}(q_{1}^{10}-1)/(q_{1}^{2}-1)=q_{1}^{2}(q^{12}-1)/(q_{1}^{2}-1).$
It follows that $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{1}\ell_{3})=1,$ contradicting Lemma 3.1$(iii)$.
$(d)$ Case $S\cong{\rm G}_{2}(p^{b}),$ where $b\geq 1.$ Here $6b=36f.$ By [5], $S$ possesses an irreducible character of degree $p^{6b}-1=q^{36}-1.$ However $q^{36}-1$ does not divide $|H|,$ a contradiction.
$(e)$ Case $S\in\{{}^{2}{\rm B}_{2}(q_{1}^{2}),{}^{2}{\rm G}_{2}(q_{1}^{2}),$ where $q_{1}^{2}=2^{2n+1},3^{2n+1},$ respectively with $n\geq 1.$ In these cases, the equation $|S|_{p}=|H|_{p}$ cannot occur.
$(f)$ Case $S\cong{}^{2}{\rm F}_{4}(q_{1}^{2}),$ where $q_{1}^{2}=2^{2n+1},$ with $n\geq 1.$ Then $12(2n+1)=36f.$ Now $\ell_{12(2n+1)}(2)\in\pi(S)\subseteq\pi(H),$ which implies that $12(2n+1)=36f\leq 18f,$ a contradiction.
$(h)$ Case $S\cong{\rm F}_{4}(p^{b})$. Then $24b=36f$ and so $2b=3f.$ By [4, 13.9], $S$ possesses a unipotent character $\varphi$ labeled by the symbol $\phi_{9,2}$ with $\varphi(1)=q_{1}^{2}\Phi_{3}^{2}(q_{1})\Phi_{6}^{2}(q_{1})\Phi_{12}(q_{1})=q^{3}\Phi_{9}^{2}\Phi_{18}.$ Hence $\varphi(1)\in{\mathrm{cd}}(L)-\{1,q^{36}\}$ and $(\varphi(1),\ell_{1}\ell_{3})=1,$ contradicting Lemma 3.1$(iii)$.
$(i)$ Case $S\cong{}^{3}{\rm D}_{4}(p^{b})$. Then $12b=36f.$ Now $\ell_{12b}(p)\in\pi(S)$ but $\ell_{12b}(p)=\ell_{36f}(p)\not\in\pi(H)$, a contradiction.
$(j)$ Case $S\cong{}^{2}{\rm E}_{6}(p^{b})$. Then $36b=36f$ and hence $b=f$ so $S\cong H.$
$(k)$ Case $S\cong{\rm E}_{7}(p^{b})$. Then $63b=36f$ and hence $7b=4f.$ By Table 1, $S$ possesses a unipotent character $\varphi\neq{\mathrm{St}}_{S}$ with $\varphi(1)_{p}=p^{46b}.$ By Lemma 3.1$(vi)$, $46b\leq 25f,$ which implies that $184b=4\cdot 46b\leq 100f=175b,$ which is impossible.
$(l)$ Case $S\cong{\rm E}_{8}(p^{b})$. Then $120b=36f$ and hence $10b=3f.$ By Table 1 and Lemma 3.1$(vi)$, $91b\leq 25f.$ It follows that $273b\leq 25\cdot 3f=250b,$ a contradiction.
$(m)$ Case $S\cong{\rm E}_{6}(p^{b})$. Then $36b=36f$ and hence $b=f.$ By [4, 13.9], $S$ has a unipotent character $\varphi$ labeled by the symbol $\phi_{6,1}$ with degree $q\Phi_{8}\Phi_{9}.$ Now $\varphi(1)\in{\mathrm{cd}}(L)-\{1,|H|_{p}\}$ is coprime to $\ell_{1}\ell_{2}\ell_{4},$ contradicting Lemma 3.1$(i)$.
The arguments for the remaining simple groups ${\rm E}_{6}(q),{\rm E}_{7}(q)$ and ${\rm E}_{8}(q)$ are similar, so we skip the details. ∎
6. Verifying Step 3: $G^{\prime}$ is isomorphic to $H$
We prove the following result in this section.
Proposition 6.1.
Let $G$ be a finite group and let $H$ be one of the finite simple groups in (1). Let $M\leq G^{\prime}$ be a normal subgroup of $G$ such that $G^{\prime}/M$ is a chief factor of $G$. Then $G^{\prime}\cong H.$
Proof.
By Propositions 4.1 and 5.1, $G^{\prime}$ is perfect and $G^{\prime}/M\cong H.$ It suffices to show that $M=1.$ By way of contradiction, assume that $M$ is nontrivial. Let $K\leq M$ be a normal subgroup of $G^{\prime}$ such that $M/K$ is a chief factor of $G^{\prime}.$ Assume first that $M/K$ is nonabelian. Then $M/K\cong S^{k},$ where $k\geq 1$ and $S$ is a nonabelian finite simple group. By Lemmas 2.1 and 2.2, $M/K$ possesses a nontrivial irreducible character $\varphi\in{\mathrm{Irr}}(M/K)$ which extends to $\varphi_{0}\in{\mathrm{Irr}}(G^{\prime}).$ As $G^{\prime}/M\cong H,$ it possesses an irreducible character $\chi$ of degree $|H|_{p},$ and hence by Gallagher’s Theorem, $\varphi_{0}\chi$ is an irreducible character of $G^{\prime}$ of degree $\varphi_{0}(1)|H|_{p}>|H|_{p}.$ As $G^{\prime}\unlhd G,$ $\varphi_{0}(1)|H|_{p}$ divides some degree of $G,$ contradicting [29, Lemma 2.4]. Thus $M/K$ is a minimal normal elementary abelian subgroup of $G^{\prime}/K.$ Let $C/K$ be the centralizer of $M/K$ in $G^{\prime}/K.$ As $M/K$ is a minimal normal elementary abelian subgroup of $G^{\prime}/K,$ we deduce that $M/K\leq C/K\unlhd G^{\prime}/K.$ In particular, $M\unlhd C\unlhd G^{\prime}.$ As $G^{\prime}/M$ is nonabelian simple, either $C=G^{\prime}$ or $C=M.$
Assume $C=M.$ Then $M/K$ is a self-centralizing minimal normal elementary abelian subgroup of $G^{\prime}/K$ and $G^{\prime}/M\cong H.$ Let $\chi\in{\mathrm{Irr}}(G^{\prime}/K)$ be such that $\chi(1)$ is the largest degree of $G^{\prime}/K.$ Applying Lemma 2.5 to $M/K\unlhd G^{\prime}/K,$ we obtain $\chi(1)>b(H).$ Now $\chi(1)\in{\mathrm{cd}}(G^{\prime})$ divides some degree of $H$ and thus $\chi(1)\leq b(H)$, a contradiction.
Assume $C=G^{\prime}$. Then $M/K=Z(G^{\prime}/K).$ Since $G^{\prime}$ is perfect, $G^{\prime}/K$ is perfect and so $M/K=(G^{\prime}/K)^{\prime}\cap Z(G^{\prime}/K),$ which implies that $G^{\prime}/K$ is a central extension of $G^{\prime}/M$ (see [13, p. 629]). We now consider each case separately.
Subcase $H\cong\textrm{F}_{4}(q),(q>2)$ or ${\rm E}_{8}(q)$. By [18, Theorem 5.1.4], the Schur multiplier of $G^{\prime}/M\cong H$ is trivial, so $G^{\prime}/K\cong G^{\prime}/M,$ which implies that $M/K$ is trivial, a contradiction.
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ By [18, Theorem 5.1.4], the Schur multiplier of $G^{\prime}/M\cong H$ is cyclic of order $d=(3,q+1).$ If $d=1,$ then $G^{\prime}/K\cong H$ so $M/K$ is trivial, a contradiction. Hence $d=3.$ Assume first that $q>2.$ Then ${}^{2}{\rm E}_{6}(q)_{sc}$ is the universal covering group of $H$ so $G^{\prime}/K\cong{}^{2}{\rm E}_{6}(q)_{sc}.$ As $G^{\prime}\unlhd G$ and ${\mathrm{cd}}(G^{\prime}/K)\subseteq{\mathrm{cd}}(G^{\prime}),$ every degree of $G^{\prime}/K$ divides some degree of $G,$ which is impossible by Lemma 3.1$(x)$. Now assume that $q=2.$ The Schur multiplier of ${}^{2}{\rm E}_{6}(2)$ is isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{6}.$ If $G^{\prime}/K$ is isomorphic to $2\cdot{}^{2}{\rm E}_{6}(2)$ or $6\cdot{}^{2}{\rm E}_{6}(2),$ then $G^{\prime}/K$ possesses an irreducible character of degree $2^{7}\cdot 3^{5}\cdot 5^{2}\cdot 7^{2}\cdot 11\cdot 13\cdot 19$ which divides no degree of ${}^{2}{\rm E}_{6}(2)$. If $G^{\prime}/K\cong 3\cdot{}^{2}{\rm E}_{6}(2),$ then $G^{\prime}/K$ has an irreducible character of degree $2^{11}\cdot 3^{3}\cdot 5\cdot 7^{2}\cdot 13\cdot 17\cdot 19$ Thus this case cannot occur.
The remaining cases can be argued similarly, we skip the details. ∎
7. Verifying Step 4: $G=G^{\prime}\times C_{G}(G^{\prime})$
In this section, we complete the proof of the main theorem. We first recall the setup in Section 3. Let $\mathcal{G}$ be a simple simply connected algebraic group in characteristic $p$ and $F$ is a Steinberg endomorphism such that $\mathcal{G}^{F}\cong H_{sc}=L,$ where $H\cong L/Z(L)$ is one of the simple groups in (1). Denote by $\mathcal{G}^{*}$ the dual group of $\mathcal{G}$ and let $F^{*}$ be the corresponding Frobenius automorphism. Then $H^{*}:=(\mathcal{G}^{*})^{F^{*}}\cong H_{ad}.$
Proposition 7.1.
Let $G$ be a finite group and let $H$ be one of the finite simple groups in (1). Suppose that ${\mathrm{cd}}(G)={\mathrm{cd}}(H).$ Then $G\cong G^{\prime}\times C_{G}(G^{\prime}).$
Proof.
By Proposition 6.1, $G^{\prime}\cong H$. Let $C=C_{G}(G^{\prime}).$ Then $G/C$ is almost simple with socle $H$ and $G^{\prime}\cap C=1$. If $G=G^{\prime}C$, then $G\cong G^{\prime}\times C$ and we are done. So, assume by contradiction that $G^{\prime}\times C<G.$ Then $G$ induces some outer automorphism, say $\sigma$, on $H.$
Subcase $H\cong\textrm{F}_{4}(q),q>2.$ In this case, $H_{sc}\cong H\cong H_{ad}.$ When $p$ is odd, $\textrm{Out}(H)$ is a cyclic group of order $f,$ consisting of field automorphisms, and when $p$ is even, $\textrm{Out}(H)$ is a cyclic group of order $2f$ with generator $\gamma$ such that $\gamma^{2}$ is a generator for the group of field automorphisms of $H.$
Assume that $\sigma$ is a nontrivial field automorphism of $H.$ Let $\ell=\ell_{12f}(p)$ be a primitive prime divisor of $p^{12f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(\mathcal{G}^{*})^{F^{*}}\cong H.$ Now $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{12}(q)$ and $H$ possesses an irreducible character $\chi_{s}$ with degree
$|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{3}^{2}\Phi_{4}^{2}\Phi_{6}^{2}\Phi_{8}.$
Let $\sigma^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\sigma.$ Then $(\mathcal{G}^{*})^{\sigma^{*}}\cong{\rm F}_{4}(p^{b}),$ where $1\leq b<f,$ so $\ell=\ell_{12f}(p)$ does not divide $|(\mathcal{G}^{*})^{\sigma^{*}}|$. By [8, Lemma 2.5(ii)], $\chi_{s}$ is not $\sigma$-invariant, and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ contradicting Lemma 3.2$(v)$.
Now assume that $\sigma$ is not a field automorphism of $H.$ Then $p=2$ and $\sigma=\gamma^{k}$ for some odd integer $1\leq k\leq f,$ where $\gamma$ is a generator of $\mathop{\rm Out}\nolimits(H).$ Let $\tau=\sigma^{2}=\gamma^{2k}.$ If $\tau$ is nontrivial, then it is a field automorphism of $H$. By applying the same argument as above, we obtain that $\chi_{s}$ is not $\tau$-invariant and so obviously $\chi_{s}$ is not $\sigma$-invariant, a contradiction. Hence $\sigma^{2}=1,$ which forces $f=k$ is odd. Thus $\sigma$ is a graph automorphism of order $2$ of ${\rm F}_{4}(2^{f}),$ where $f\geq 2$ is odd. By [21, Theorem 2.5(e)], $H$ has two unipotent characters labeled by the symbol $\phi_{2,16}^{\prime}$ and $\phi_{2,16}^{\prime\prime}$ with degree $\frac{1}{2}q^{13}\Phi_{4}\Phi_{8}\Phi_{12},$ which are fused under $\sigma.$ Therefore $G/C$ possesses an irreducible character of degree $q^{13}\Phi_{4}\Phi_{8}\Phi_{12}.$ However this contradicts Lemma 3.2$(vi)$ as $q$ is even.
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ Note that $\textrm{Out}(H)$ is a group of order $2df,$ where $d=(3,q+1).$ Assume first that $G/C$ possesses only inner and diagonal automorphism. It follows that $G/C\cong{}^{2}{\rm E}_{6}(q)_{ad},$ $3\mid(q+1)$ and $\sigma$ is a diagonal automorphism of $H.$ Thus ${\mathrm{cd}}({}^{2}{\rm E}_{6}(q)_{ad})={\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G)={\mathrm{cd}}(H),$ which is impossible by Lemma 3.1$(x)$.
Hence we can assume that $\sigma=\gamma\varrho,$ where $\varrho$ is a field automorphism and $\gamma$ is either trivial or a diagonal automorphism. Let $\ell=\ell_{18f}(p)$ be a primitive prime divisor of $p^{18f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(\mathcal{G}^{*})^{F^{*}}.$ Now $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{18}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree
$|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{}^{2}{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{18}$
which acts trivially on its center, so $\chi_{s}\in{\mathrm{Irr}}(H).$ Let $\varrho^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\varrho.$ Then $(\mathcal{G}^{*})^{\varrho^{*}}\cong{}^{2}{\rm E}_{6}(p^{b})_{ad},$ where $1\leq b<f,$ so $\ell=\ell_{18f}(p)$ does not divide $|(\mathcal{G}^{*})^{\varrho^{*}}|,$ and hence by [8, Lemma 2.5(ii)], $\chi_{s}$ is not $\varrho$-invariant. As $\chi_{s}$ is invariant under the diagonal automorphisms, $\chi_{s}$ is not $\sigma$-invariant and so some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible.
Subcase $H\cong\textrm{E}_{6}(q).$ Note that $\textrm{Out}(H)\cong\langle\gamma\rangle(\langle\varrho\rangle\times\langle\tau\rangle))$ is a group of order $2df,$ where $d=(3,q-1),$ and $\gamma$ is a diagonal automorphism of order $d,$ and $\varrho$ is a field automorphism of order $f$ and $\tau$ is a graph automorphism of order $2$ (see for example [10, Theorem 2.5.12]). Moreover $\tau$ inverts the cyclic group $\langle\gamma\rangle.$
Assume first that $G/C$ possesses only inner and diagonal automorphism. It follows that $G/C\cong{\rm E}_{6}(q)_{ad},$ $3\mid(q-1)$ and $\sigma$ is a diagonal automorphism of $H.$ Thus ${\mathrm{cd}}({\rm E}_{6}(q)_{ad})={\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G)={\mathrm{cd}}(H),$ which is impossible by Lemma 3.3$(x)$.
Assume next that $\sigma=\gamma\varrho,$ where $\varrho$ is a nontrivial field automorphism and $\gamma$ is either trivial or a diagonal automorphism. Let $\ell=\ell_{12f}(p)$ be a primitive prime divisor of $p^{12f}-1$ and let $s$ be a semisimple element of order $\ell$ in $K\cong{\rm F}_{4}(q)$ where $K\leq(H^{*})^{\prime}$ is chosen so that it is fixed under the involutory graph $\tau$ (see [10, Proposition 4.9.2]). Now $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{12}\Phi_{3}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree
$|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{12}\Phi_{3}$
which is trivial on the center as $s\in(H^{*})^{\prime}$ so $\chi_{s}\in{\mathrm{Irr}}(H).$ Let $\varrho^{*}$ be an automorphism of $\mathcal{G}^{*}$ induced from the field automorphism $\varrho.$ Then $(\mathcal{G}^{*})^{\varrho^{*}}\cong{\rm E}_{6}(p^{b})_{ad},$ where $1\leq b<f,$ so $\ell=\ell_{12f}(p)$ does not divide $|(\mathcal{G}^{*})^{\varrho^{*}}|$. By [8, Lemma 2.5(ii)], $\chi_{s}$ is not $\varrho$-invariant. As $\chi_{s}$ is invariant under the diagonal automorphisms, it is not $\varrho$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible.
Now assume that $\sigma=\gamma\varrho\tau,$ where $\gamma$ is either trivial or a diagonal automorphism, $\varrho$ is a nontrivial field automorphism and $\tau$ is a graph automorphism. As the field and graph automorphisms commute by [10, Theorem 2.5.12], we can write $\sigma=\gamma\tau\varrho.$ Choose $s\in(H^{*})^{\prime}$ as in previous paragraph with $|s|=\ell_{12}(q).$ Since $s$ in $\tau$-invariant, [26, Corollary 2.5] yields that $\chi_{s}^{\tau}=\chi_{\tau(s)}=\chi_{s}.$ As $\chi_{s}$ is also $\gamma$-invariant, by applying [26, Corollary 2.5] again $\chi_{s}$ is $\gamma\tau$-invariant. Thus $\chi_{s}^{\sigma}=\chi_{s}^{\gamma\tau\varrho}=\chi_{s}^{\varrho}\neq\chi_{s}.$ Therefore $\chi_{s}$ is not $\sigma$-invariant, which leads to a contradiction.
Finally assume that $\sigma=\gamma\tau,$ where $\gamma$ is either trivial or a diagonal automorphism and $\tau$ is a graph automorphism. Let $\ell=\ell_{9f}(p)$ be a primitive prime divisor of $p^{9f}-1$ and let $s$ be a semisimple element of order $\ell$ in $(H^{*})^{\prime}.$ Let $T$ be a maximal torus of $H^{*}$ containing $s.$ It is well known that $T$ is uniquely determined up to conjugation in $H^{*}.$ Now $|C_{(\mathcal{G}^{*})^{F^{*}}}(s)|=\Phi_{9}$ and $\mathcal{G}^{F}$ possesses an irreducible character $\chi_{s}$ with degree
$|(\mathcal{G}^{*})^{F^{*}}:C_{(\mathcal{G}^{*})^{F^{*}}}(s)|_{p^{\prime}}=|{\rm E}_{6}(q)_{ad}|_{p^{\prime}}/\Phi_{9},$
which is trivial on the center as $s\in(H^{*})^{\prime}$ so $\chi_{s}\in{\mathrm{Irr}}(H).$ As $\chi_{s}$ is invariant under the diagonal automorphisms, it suffices to show that $\chi_{s}$ is not $\tau$-invariant: if so, then it is not $\sigma$-invariant and hence some proper multiple of $\chi_{s}(1)$ is a degree of $H,$ which is impossible. We know that ${\mathbf{C}}_{H^{*}}(\tau)\cong\textrm{F}_{4}(q)$ and thus $\ell\nmid|{\mathbf{C}}_{H^{*}}(\tau)|.$ Consider the $H^{*}$-conjugacy class $(s).$ Since $|s|=\ell,$ $H^{*}$-conjugacy class $(s)$ is not $\tau$-invariant. By [26, Corollary 2.5], $\chi_{s}$ is not $\tau$-invariant.
The arguments for the remaining cases are similar. ∎
We now complete the proof of the main theorem.
Proof of Theorem 1.1.
Let $G$ be a finite group and let $H$ be a finite simple exceptional group of Lie type. In view of [12, 29, 30, 32, 33, 34], we assume that $H$ is one of the simple groups listed in (1). By Propositions 6.1 and 7.1, $G^{\prime}\cong H$ and $G\cong G^{\prime}\times C_{G}(G^{\prime})$. It follows that $C_{G}(G^{\prime})\cong G/G^{\prime}$ is abelian. Hence $G\cong G^{\prime}\times A$, where $A=C_{G}(G^{\prime})$ is an abelian group. This completes the proof of the theorem. ∎
Acknowledgment
The author is grateful to the reviewer and the editor for numerous suggestions and corrections that improve the exposition of the paper. The author also thanks Frank Lübeck for his help during the preparation of this paper.